### v2

# Quantum Vacuum energy as the origin of Gravity

Abstract
[…]
I Introduction
Gravity remains the least understood of the fundamental forces, in spite of the discoveries of General Relativity (GR) and Quantum Mechanics (QM) over 100 years ago, this year being recognized as the 100-th anniversary of a complete theory of QM due to Schrödinger and Heisenberg, although the birth of QM goes back to Planck’s black body studies at finite temperature. Classical singularities at the center of black holes and at the time of the Big Bang have not yet been understood and resolved in classical GR. Many physicists expect that such singularities will be resolved by a theory of quantum gravity, but it’s fair to say that a complete theory of quantum gravity has not been forthcoming, although string theory offers some promise. For these reasons it is worthwhile to pursue new ideas on the origin or emergence of Gravity itself, especially if such ideas suggest that perhaps Gravity doesn’t need to be quantized in the first place. Such a theory should be able to explain the value of Newton’s constant $G_{N}$ itself.
In this article we explore the question “Can Quantum vacuum energy be the origin of Gravity?” [^2] This article was initially intended simply as a theoretical exercise in the form of an essay to address this question, but we found it can lead to a fruitful new perspective on Gravity more generally, and it is not without some predictions, if the basic ideas continue to hold. This perspective is based on the idea that perhaps we can learn more about the foundations of Gravity by considering the cosmological far future, in contrast to the far past where much less is understood, in particular the origin of the Big Bang and possibly inflation. The late Universe is dominated by so-called Dark Energy, thus one should first attempt to understand the full implications of this in a Universe where Dark Energy is its sole component. By assumption, we will equate Dark Energy with the quantum vacuum energy density $\rho_{\rm vac}$} computed in quantum field theory (QFT) in flat local Minkowski space, defined as
$\langle 0|T_{\mu\nu}|0\rangle={\rho_{\rm vac}}\,\eta_{\mu\nu}$ (1)
with the convention $\{\eta_{\mu\nu}\}={\rm diag}\{1,-1,-1,-1\}$ where the Vacuum $|0\rangle$ is the ground state for all quantized fields, excluding gravity. The gravitational implications of the quantum field theoretic properties of $\rho_{\rm vac}$} will be then explored.
The first observations of Gravity were of course at the surface of the Earth which is manifested as a constant acceleration $g$ for all matter regardless of mass or specific material. Newton later understood $g$ to be “environmental” rather than fundamental, namely $g=G_{N}M_{E}/R_{E}^{2}$, where $M_{E},R_{E}$ are the mass and radius of the Earth. In Newton’s universal theory of Gravity, $G_{N}$ is considered a new fundamental constant of Nature, and this view was adopted by Einstein in his formulation of General Relativity. [^3] Throughout this paper, $G_{N}=6.67\cdot 10^{-11}{\rm m^{3}}/{\rm kg\,s^{2}}$ is Newton’s constant as measured today. We keep the fundamental constants $\hbar$ and $c$ explicit in order to make certain conceptual points and to provide some numbers with units. “Natural” units based on the Planck length $\ell_{p}$ can be deceptive and are not so natural for this article. In contrast, below we will argue that $G_{N}$ is also environmental, namely
$G_{N}=\dfrac{c^{4}}{2}\dfrac{{{\rm R}_{\scriptscriptstyle\infty}}}{{{\rm E}_{\scriptscriptstyle\infty}}}=\dfrac{c^{2}}{2}\dfrac{{{\rm R}_{\scriptscriptstyle\infty}}}{{{\rm M}_{\scriptscriptstyle\infty}}}$ (2)
[…]
Based on the above remarks, what will be central to our exercise is the quantity $\rho_{\rm vac}$}. For purposes of illustration, consider a Universe which only consists of a single harmonic oscillator with ground state energy $E_{0}=\hbar\omega/2$, in addition to an imagined observer. Since classical mechanics in unaffected by shifts of the potential energy by an arbitrary constant, $E_{0}$ is not measurable in classical physics. If all that exists is the harmonic oscillator in empty space, then $E_{0}$ also cannot be measured in a single experiment based on QM. There are two known ways to measure $E_{0}$. First, one can couple the harmonic oscillator to a generic heat bath at temperature $T$. Then one can in principle infer $E_{0}$ from the high temperature behavior of the partition function or its logarithm which is the free energy. It’s important that this is based on the free energy and not its derivatives, and thus it is not invariant under shifts of the potential energy. The second manner to measure $E_{0}$ is to couple the oscillator to Gravity, since gravity originates from all the matter-energy in the Universe and is sensitive to its ground state energy. The fact that the ground state energy can be measured either by coupling to a heat bath or gravity suggests parallels between thermodynamics and gravity that have already been recognized. Deceptively simple as it may appear, to study QFT at finite temperature with periodic boundary conditions, one necessarily needs to study quantum fields in curved Euclidean space. In 1 spatial dimension the Euclidean space is a torus, and the problem of conformal theory on a torus is a completely solved problem. Like Gravity, the Vacuum $|0\rangle$ is omnipresent. The Vacuum “knows” the physical properties, such as mass, quantum numbers, and interactions of any particle excitations above it, and is perhaps the only entity that exists everywhere in the Universe and for all times. Everything else is some quantum excited state, including people, and typically unstable. Nature does not abhor a Vacuum, rather it depends on it. [^5] To paraphrase with translation a recognizable philosopher, If you look deeply enough into the Void, it eventually looks back at you. So it is reasonable to suppose that the Vacuum, in particular its ground state energy density $\rho_{\rm vac}$}, is at the origin of gravity. In other words, if the only way to measure $E_{0}$ is with gravity, then reversing the argument, perhaps quantum vacuum energy is the origin of Gravity. For the same reason that one does not necessarily have to quantize the heat bath to study finite temperature QFT, perhaps one does not have to quantize Gravity, however this deep issue will be left as an open question. This does not preclude the existence of classical gravitational waves that have been measured, as they can be viewed as classical perturbations of the Vacuum as an elastic medium. This article explores this possibility as a theoretical academic exercise to see where it leads.
This article assumes that $\rho_{\rm vac}$} is well-defined in Minkowski space, namely it is as well defined as the ground state energy of a harmonic oscillator. Early considerations of $\rho_{\rm vac}$} based on viewing a free quantized field as a collection of harmonic oscillators with energies $\hbar\omega_{\bf{k}}$, with $\omega_{\bf{k}}=\sqrt{{\bf{k}}^{2}+m^{2}}$ led to the introduction of a cut-off in $|{\bf{k}}|$. If this cut-off is taken to be based on the Planck length $\ell_{p}\equiv\sqrt{\hbar G_{N}/c^{3}}$ then $\rho_{\rm vac}$} is much too high by even 100 orders of magnitude and this is perhaps the first formulation of the “cosmological constant problem” Weinberg . This naive computation of $\rho_{\rm vac}$} depends on the cut-off and goes to infinity as the cut-off goes to infinity. This computation is essentially meaningless, as is the often quoted “worse prediction in physics by $10^{120}$ orders of magnitude”, simply because it is not a real prediction. In quantum field theory one is accustomed to dealing with ultra-violet (UV) singularities. In recent work ALrhovac1 ; ALrhovac2 we formulated a proper computation of $\rho_{\rm vac}$} which renders it finite and well-defined. Since this will be central to subsequent arguments we list its main properties.
(i)   For 2 dimensional integrable theories $\rho_{\rm vac}$} is exactly computable from the Thermodynamic Bethe Ansatz, and also from the S-matrix bootstrap, and takes the form ${\rho_{\rm vac}}=m_{1}^{2}/2\mathfrak{g}$ where $m_{1}$ is the lightest mass particle. In the works ALrhovac1 ; ALrhovac2 this formula was extended to 4 spacetime dimensions:
${\rho_{\rm vac}}=\dfrac{3}{4}\dfrac{c^{5}}{\hbar^{3}}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}}$ (3)
where $m_{\rm z}$ is the physical mass of the lightest particle, which we will refer to as the “zeron”, where zero signifies the Vacuum, and $\mathfrak{g}$ is a dimensionless interaction coupling constant. Although one does not have the powerful simplifications of integrability in 4 dimensions, this formula was well motivated in ALrhovac1 ; ALrhovac2 and should not be viewed as an ad hoc assumption, although unlike in 2 dimensions it hasn’t yet been fully rigorously proven since the $1/\mathfrak{g}$ dependence indicates it is highly non-perturbative. Clearly the above expression is purely quantum mechanical as can be seen from its $\hbar$ dependence. The $1/\mathfrak{g}$ dependence can be anticipated, and is in fact necessary, since $\rho_{\rm vac}$} should diverge as $\mathfrak{g}\to 0$ due to an unavoidable logarithmic divergence for free QFT in the UV when the spatial dimension $d$ is evenALrhovac1 . [^6] This is essentially another, but refined, version of the “Cosmological Constant Problem”. The precise identification of the zeron as a real particle will not be necessary for our arguments, thus this identification will be left as an open question. However it was argued in ALrhovac2 that $m_{\rm z}$ is fixed in the deep UV above the electro-weak spontaneous symmetry breaking scale, and that Majorana neutrinos are viable candidates (see (8)). [^7] For experimental constraints on neutrino masses see for instance NeutrinoMasses .
[…]
${\cal F}=-\dfrac{\pi^{2}T^{4}}{90}\,c(r),~~~~~~{\rm with}~~c(r)=c_{\rm uv}+c_{4}\,r^{4}+\ldots$ (4)
for some constant coefficients $c_{\rm uv},c_{4}$. [^8] In 2 spacetime dimensions $c_{\rm uv}$ is the effective Virasoro central charge. Then ${\rho_{\rm vac}}=-c_{4}\pi^{2}m_{\rm z}^{4}/90$. This was used in ALrhovac1 to estimate $\rho_{\rm vac}$} for QCD based on finite temperature lattice results.
(iii)   $\rho_{\rm vac}$} can be independently computed from the form-factor bootstrap that relates the two particle form-factor to the zero particle one, which is a 1-point vacuum expectation value of the field in question ALrhovac2 . This is the reason for $m_{\rm z}$ to be the lightest mass particle, since in principle the S-matrix and form factors for other particles can be obtained by bootstrapping the lightest mass particle. Applying this to $\langle 0|T_{\mu\nu}|0\rangle$, this leads to a determination of the coupling $\mathfrak{g}$. As usual, defining $\cal T$} from the S-matrix $S=1+i{\cal T}$, then
$\lim_{s\to\infty}{\cal T}(s)=\dfrac{\mathfrak{g}}{m_{\rm z}^{2}}\dfrac{1}{s},$ (5)
where $s=(p_{1}+p_{2})^{2}$ is the Mandelstam variable for 2-particle scattering.
In two spacetime dimensions, the two above definitions of $\rho_{\rm vac}$} were shown to agree exactly for a variety of integrable quantum field theories ALrhovac1 ; ALrhovac2 . As previously mentioned, in one spatial dimension it takes the form ${\rho_{\rm vac}}=c^{3}m_{1}^{2}/2\hbar\mathfrak{g}$ where $m_{1}$ is the lightest particle and $\mathfrak{g}$ can be explicitly computed. Various consistency checks on this formula were made, including that it must vanish for supersymmetric theories and also theories with a fractional supersymmetries where the hamiltonian is in the center of the universal enveloping algebra of the conserved charges. (See ALrhovac1 and references therein.) The coupling $\mathfrak{g}$ can be positive or negative, and $\rho_{\rm vac}$} can even oscillate in sign for a given model, for instance for the sine-Gordon model where it oscillates from positive to negative, where it is zero at the fractional supersymmetric points. In 4 spacetime dimensions it was not rigorously shown that equations (4), (5) lead to the same formula for $\rho_{\rm vac}$}, thus (3) should still be viewed as well-motivated but still conjectural. For the remainder of the article we simply assume the well-motivated formula (3).
Some independent support for the result (3) comes from Swampland ideas Festina1 ; Festina2 . By studying a charged particle of mass $m$ in the presence of a black hole, it was argued that
${\rho_{\rm vac}}<\dfrac{m^{4}}{2e^{2}}$ (6)
where here $\mathfrak{g}$ is identified with the quantum electrodynamic (QED) fine structure constant $\alpha=e^{2}/\hbar c$. Actual QED will not play a role in this article since we assume the zeron is electrically neutral, however, as we will see, it will be instructive to compare couplings and renormalization group parameters with QED, just to check if they are reasonable. Our result (3) is stronger than (6) since it is not an inequality, but it is still consistent with (6).
As previously stated, in this article we do not propose what the hypothetical zeron particle actually is, nor even that it is a detectable particle, however we can constrain it’s mass $m_{\rm z}$ from the cosmologically measured value of $\rho_{\rm vac}$}. Taking it to be ${\rho_{\rm vac}}=\Omega_{\Lambda}\rho_{\rm crit}$, where $\Omega_{\Lambda}\approx 0.68$ one finds Planck
${\rho_{\rm vac}}=5.2\times 10^{-10}\,\dfrac{{\rm Joule}}{{\rm m}^{3}}.$ (7)
Using this value, then (3) implies
$m_{\rm z}=0.0024\,\mathfrak{g}^{1/4}\,{\rm eV}.$ (8)
For $\mathfrak{g}={\cal O}(1)$, it is interesting to note that $m_{\rm z}$ is on the order of proposed neutrino masses NeutrinoMasses . Below we will propose additional constraints on the coupling $\mathfrak{g}$.
It is instructive to compare the perspective on Gravity developed in this article with some proposed theories of quantum gravity, since this clearly delineates some important distinctions. We take “quantum gravity” to mean a theory where the space-time metric $g_{\mu\nu}(x)$ itself is quantized which implies the existence of spin 2 gravitons. Although the definition of the Planck length $\ell_{p}$, defined from $G_{N}$ as $G_{N}=c^{3}\,\ell_{p}^{2}/\hbar$, is a priori just based on dimensional analysis, one expects that $\ell_{p}$ is a relevant scale when considering graviton-graviton scattering. The latter does not play a role in this article, which makes one question whether $\ell_{p}$ is a fundamental constant, however this issue will not be completely settled in the present cosmological context. [^9] Based on this article, units based on $\hbar$, $c$ and $m_{\rm z}$ are more natural. A significant result of string theory is that it provides a consistent prescription for the computation of graviton scattering that is finite StringTheory . Thus gravity “emerges” from string theory, in the sense that Newton’s constant $G_{N}$ is determined by matching graviton scattering with low energy gravity. This leads to a formula for $G_{N}$:
$G_{N}\sim c_{Y}\,g_{s}^{2}\,\,\dfrac{c^{3}\ell_{s}^{2}}{\hbar}$ (9)
[…]
With the above introductory remarks, let us summarize the remainder of this article. The two primary assumptions we make are the formula (3) for $\rho_{\rm vac}$} which was well-motivated by the works ALrhovac1 ; ALrhovac2 , and the renormalization group flow for $\mathfrak{g}$ in equation (35) below. In the next section we formulate a gravitational version of the electromagnetic Casimir effect, and provide an argument for how gravity can arise from $\rho_{\rm vac}$}. In particular we show how Einstein’s field equations emerge, in this context in the form of Friedmann’s equations. The formula (2) leads to new interpretation of the Gibbons-Hawking entropy of de Sitter space, wherein the quantum informational “bits” are identified as quantized massless particles at the horizon with wavelength $\lambda=2\pi{{\rm R}_{\scriptscriptstyle\infty}}$. The analysis in Section II is independent of the form $\rho_{\rm vac}$} in (3), and we turn to understanding the implications of (3) in Section III. There we assume $\mathfrak{g}$ is a marginal coupling with a renormalization group (RG) flow, which as we will argue is well-motivated, and show that this leads to an effective, induced RG flow for Newton’s constant $G_{N}$ as a function of an energy scale $\mu$. [^11] We previously published some preliminary work on the possible implications of the RG flow of $\mathfrak{g}$ for cosmology in LeClairUniverse , however the present article supersedes it. As is standard in Cosmology, we tie this energy scale $\mu$ to the redshift $z$ and obtain equation (43). Solving the resulting equations for the scale factor $a(t)$ results in a significant modification of de Sitter space that is essentially an inverted gaussian and at no time is $a(t)=0$, which signifies no classical type of geometric singularity since for all times $a(t)>a_{\rm min}$. The solution to the scale factor satisfies an interesting symmetry between the far past and far future, namely $a(t)=a(-t+2t_{\rm min})$ where $a(t_{\rm min})=a_{\rm min}$. There is no Big Bang in our model since the solutions in past and future match at the time $t_{\rm min}$. Extending our solution to the far past, the time evolution of the Universe is more analogous to the swing of a pendulum. Our interpretation of this is that the age of the Universe is actually infinite since the solution to $a(t)$ extends to the far past without ever reaching the singularity at $a=0$. In Section IV we add matter and radiation, and due to the RG flow of $G_{N}$ in (43) there is again no $a=0$ singularity and the symmetry between the past and future continues to hold. This implies very small logarithmic corrections to the $\Lambda{\rm CDM}$ model if $\widehat{b}$ is small. We apply these ideas to the so-called Hubble tension, wherein the discrepancy of measurements of Hubble constant $H_{0}$ is proposed to arise from the fact that the two contradictory measurements rely on data that refers to different epochs where Newton’s constant is effectively different since the energy scales involved are different.
II Emergence of Einstein’s equations from a Gravitational Casimir effect
In this section we formulate a gravitational version of the electromagnetic Casimir effect. Without assuming Einstein’s field equations, we present heuristic arguments for how they emerge in this context. These arguments lead to the formula (2) for Newton’s constant $G_{N}$ expressed in terms of length and energy scales in the Universe. This doesn’t rely on the explicit formula (3) for $\rho_{\rm vac}$}, where its implications will be studied in the next section.
Let us first review the electromagnetic Casimir effect, which is well understood theoretically and has been measured Casimir . This Casimir effect requires two uncharged, parallel conducting plates in vacuum separated by a distance $L$. The energy density of quantized photons between the plates $\rho_{\rm rad}$ leads to a measurable force between the plates $F=\partial_{L}E$ where $E=\rho_{\rm rad}\cdot V$ where $V$ is the volume between the plates. It is important to point out that the theoretical predictions, which are confirmed in the measurements, do not measure the vacuum energy density $\rho_{\rm vac}$} but rather the energy density of free photons to leading order, whose pressure is $p=w\rho_{\rm rad}$, where $w=1/3$, whereas genuine vacuum energy density has w=1w=-1. In fact, since the free theory of photons is conformally invariant, it’s vacuum energy density ${\rho_{\rm vac}}=0$.
II.1 Formulating a gravitational Casimir effect without plates
For gravity, there is no well-formulated analog of the electromagnetic Casimir effect, since hypothetical gravitons cannot be constrained by material boundaries like plates. In any case, if any experiment could be imagined that overcomes these obvious difficulties, its effects would be much too small to be measurable. This led us to formulate a gravitational Casimir effect without material boundaries, namely without plates, that is sensitive to $\rho_{\rm vac}$}, where the latter is the vacuum energy density of all quantum matter and radiation excluding gravity. Throughout this paper we assume ${\rho_{\rm vac}}>0$ such that we are dealing with de Sitter space. We make three additional assumptions: (i) $\rho_{\rm vac}$} is well-defined and finite in flat Minkowski space for the entirety of quantum fields for matter and radiation without gravity, defined in (1). This quantity $\rho_{\rm vac}$} is in principle computable without gravity, as explained in the Introduction. Secondly, (ii) The Vacuum has no time dependence originating from Einstein’s equations, namely $\partial_{t}{\rho_{\rm vac}}=0$. However as we will see in the next section, tying the RG energy scale to the redshift $z$ leads to an additional induced time dependence. Third, (iii) There is no additional and independent free parameter $\Lambda$ in Einstein’s equations corresponding to a classical cosmological constant. This is consistent with the fact that $\rho_{\rm vac}$} is not invariant under arbitrary shifts of the potential energy since it is in principle computable from the logarithm of the partition function and not its derivatives.
We carry out this construction in arbitrary spatial dimension $d$ in order to make certain points. Consider first a relativistic QFT in $d=1$ spatial dimension, from which one can compute $\rho_{\rm vac}$}. There are an infinite number of integrable QFT’s where one can calculate $\rho_{\rm vac}$} exactly; see for instance the examples in ALrhovac1 ; ALrhovac2 . Let the spatial dimension be a finite size segment of length $2R$. Since total vacuum energy is $E=2R\,{\rho_{\rm vac}}$, there is a force $F=dE/dR=2{\rho_{\rm vac}}$ that can in principle be measured by moving the “walls”, in this case simply the endpoints of the segment. However if the endpoints do not consist of matter being held in place, the effect of this force is expansion or contraction of the size of the segment, depending on the sign of $\rho_{\rm vac}$}. This naturally leads us to introduce geometry, namely a metric $g_{\mu\nu}$:
$ds^{2}=-g_{\mu\nu}(x)dx^{\mu}dx^{\nu}=-c^{2}dt^{2}+a(t)^{2}dx^{2}.$ (10)
The dimensionless scale factor can be thought of as $a(t)=R(t)/R_{0}$ where $R_{0}$ is the scale factor today, such that at the present time $t_{0}$, $a(t_{0})=1$. In order to make sense of Newton’s 2nd law, one needs to introduce an inertial mass $M$. One also needs to introduce a length scale $R$ such that $R\ddot{a}$ is an acceleration, where as usual over-dots refer to time derivatives. Equating the force $F$ with $2{\rho_{\rm vac}}$, and requiring both sides scale the same way with the scale factor $a$, one obtains
$MR\,\ddot{a}=2a\,{\rho_{\rm vac}}~~~\Longrightarrow~~\dfrac{\ddot{a}}{a}=\dfrac{2{\rho_{\rm vac}}}{MR}.$ (11)
We next require $\partial_{t}{\rho_{\rm vac}}=0$. One needs another equation to enforce this, which is $\ddot{a}/a=(\dot{a}/a)^{2}$:
$\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{2{\rho_{\rm vac}}}{MR}.$ (12)
One can easily check that one time derivative of (12) combined with (11) implies $\partial_{t}{\rho_{\rm vac}}=0$.
Let us extend the above arguments to $d$ spatial dimensions, where now $R$ is a radius. Define $V_{d}(R)$ to be the volume of a sphere in $d$ dimensions. Then one has
$MR\,\ddot{a}=\dfrac{a}{d}\,\dfrac{\partial V_{d}(R)}{\partial R}{\rho_{\rm vac}},~~~~~V_{d}(R)=\dfrac{2\pi^{d/2}R^{d}}{d\,\Gamma(d/2)}.$ (13)
The factor of $1/d$ on the right hand side of the first equation is due to the vectorial nature of the force. To summarize, we have two equations after imposing $\partial_{t}{\rho_{\rm vac}}=0$:
$\dfrac{\ddot{a}}{a}=\dfrac{R^{d-2}}{M}\left(\dfrac{2\pi^{d/2}}{d\,\Gamma(d/2)}\right){\rho_{\rm vac}}$ (14)
$\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{R^{d-2}}{M}\left(\dfrac{2\pi^{d/2}}{d\,\Gamma(d/2)}\right){\rho_{\rm vac}}.$ (15)
Since $\rho_{\rm vac}$} is independent of time, $H\equiv\dot{a}/a$ is a constant. If one identifies $H=c/R$, then the second equation just implies
${\rho_{\rm vac}}\cdot V_{d}(R)=Mc^{2},$ (16)
such that $Mc^{2}$ can be interpreted as the total vacuum energy in a sphere of radius $R$.
Thus far, Newton’s constant $G_{N}$ has played no role, since we have not assumed Newtonian gravity nor Einstein’s general relativity. $G_{N}$ can be identified by comparing with Einstein’s field equations for this particular metric. We adopt standard conventions for the Einstein-Hilbert action in $d$ dimensions:
${\cal S}=\dfrac{1}{16\pi G_{d}}\int d^{d+1}x\,\sqrt{-g}\left({\cal R}+{\cal L}_{\rm matter}\right).$ (17)
The metric is as in (10), with with $dx^{2}$ replaced with $d\vec{x}\cdot d\vec{x}$. There is no a priori reason to include a curvature term proportional the standard $k\in\{-1,0,1\}$ in the spatial metric based on the above formulation of the Casimir effect. In other words $k=0$ is natural for our formulation, since, from our perspective, gravity originates from $\rho_{\rm vac}$} in Minkowski space, and the latter clearly has $k=0$. This leads to the field equations
${\cal R}_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}\,{\cal R}=\dfrac{8\pi G_{d}}{c^{d+1}}\,T_{\mu\nu}.$ (18)
If one carefully keeps track of the $d$-dependence, this leads to the following Friedmann equations [^12] The $d$ dependence of this formula should be a well-known result. One can find it for instance in Gibbons .
$\dfrac{\ddot{a}}{a}=-\dfrac{8\pi G_{d}}{d(d-1)c^{2}}\left[(d-2)\rho+d\,p\right]$ (19)
$\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{16\pi G_{d}}{d(d-1)c^{2}}\,\rho.$ (20)
Specializing to vacuum energy with $T_{\mu\nu}=\langle 0|T_{\mu\nu}|0\rangle$ in (1), where $p=-{\rho_{\rm vac}}$, one obtains the equations (14) with the identification
$G_{d}=c^{2}\dfrac{R^{d-2}}{M}\left(\dfrac{(d-1)\pi^{(d-2)/2}}{8\Gamma(d/2)}\right).$ (21)
[…]
It remains to specify $R$ and $M$ in the above equations. From our formulation of the gravitational Casimir effect, it’s clear that $M$ and $R$ refer to large scale properties of the Universe, and this requires some input from the solution $a(t)$. Recall that so far we have only considered vacuum energy density in $T_{\mu\nu}$, thus we are dealing with de Sitter space where $H\equiv\dot{a}/a$ is a constant by the second Friedmann equation (20). The natural choice is then
$R=R_{H}=\dfrac{c}{H},$ (22)
which is referred as the Hubble radius. With the identification of $G_{d}$ in (21), the equation (20) is equivalent to
${\rho_{\rm vac}}V_{d}(R_{H})=Mc^{2}.$ (23)
In de Sitter space, $R_{H}$ has several equivalent meanings which are all well-known and we briefly review. Here
$a(t)=e^{H(t-t_{0})}.$ (24)
$R_{H}$ equals the event horizon, which marks the boundary beyond which an observer at a given time can never receive signals from events due to the accelerating expansion. For a co-moving observer in de Sitter space, the proper distance is given by
$d_{\rm event}=a(t)\int_{t}^{\infty}\dfrac{c\,dt^{\prime}}{a(t^{\prime})}=R_{H}.$ (25)
The Hubble volume $\tfrac{4}{3}\pi R_{H}^{3}$ represents the volume that will remain causally connected to an observer as $t\to\infty$. In other words, it is the maximal volume of space that can influence an observer’s world line. $R_{H}$ also represents the distance at which objects recede in the expansion at the speed of light, which perhaps explains why the formula (2) is the correct relation for the mass $M_{\bullet}$ and radius $R_{\bullet}$ of a black hole if one identifies $M_{\bullet}={{\rm M}_{\scriptscriptstyle\infty}},R_{\bullet}={{\rm R}_{\scriptscriptstyle\infty}}$. Based on the above discussion we identify $R=R_{H}\equiv{{\rm R}_{\scriptscriptstyle\infty}}$, and $M={{\rm M}_{\scriptscriptstyle\infty}}$ defined through (23) where the subscripts refer to the far future. In 3 spatial dimensions, this gives the formula in the Introduction:
$G_{N}=\dfrac{c^{2}}{2}\dfrac{{{\rm R}_{\scriptscriptstyle\infty}}}{{{\rm M}_{\scriptscriptstyle\infty}}}.$ (26)
Let us for the moment express (26) as $G_{N}=3c^{2}H^{2}/8\pi{\rho_{\rm vac}}$. If one relies only on experimental data in modern cosmology, then the above expression (26) for $G_{N}$ is approximately correct with some assumptions, however it should be realized that at this stage this is a tautology since analysis of the cosmological data depends on $G_{N}$. First of all, we have not yet included matter and radiation as excitations over the Vacuum, where $H$ is no longer a constant, and we will return to this below. Let us simply point out that if one takes the experimental value (7) and identifies $H=\sqrt{\Omega_{\Lambda}}H_{0}$ where $H_{0}$ is Hubble constant as measured by the CMB (52) and $\Omega_{\Lambda}=0.68$, then the above expression evaluates to $G_{N}=6.72\times 10^{-11}$ ${\rm m}^{3}/{\rm kg}\,{\rm s}^{2}$. Let us note that the same computation using $H_{0;{\rm SN}}$ in (52) yields a higher value for $G_{N}$. This will be addressed below in connection with the “Hubble tension” by taking into account the explicit expression (3) for $\rho_{\rm vac}$}.
II.2 An interpretation of the Gibbons-Hawking entropy of de Sitter space
There are some important results in semi-classical gravity where the gravitational field is not quantized, namely they do not rely on the existence of gravitons, but the result still depends on the Planck length $\ell_{p}$, such as the Bekenstein-Hawking entropy and temperature for black holes Bekenstein ; Hawking . In this present context, the analog is the Gibbons-Hawking entropy for de Sitter space GibbonsHawking . Since these results depend on $G_{N}$, the expression (2) for $G_{N}$ leads to an interesting new interpretation of this entropy.
[…]
$S=k_{B}N_{\rm bits}.$ (27)
The Gibbons-Hawking entropy was thus proposed to be
$S_{GH}\equiv\dfrac{k_{B}}{4}\dfrac{A}{\ell_{p}^{2}}=k_{B}\left(\dfrac{\pi c^{3}}{\hbar}\right)\dfrac{{{\rm R}^{2}_{\scriptscriptstyle\infty}}}{G_{N}}.$ (28)
where $A=4\pi{{\rm R}^{2}_{\scriptscriptstyle\infty}}$. Here each bit is commonly interpreted as a pixelation of the area $A$ into quantized units of area $4\ell_{p}^{2}$, suggesting that spacetime itself is somehow quantized into these minimal area units at the horizon. Such a quantization of spacetime itself remains a somewhat vague notion, and difficult to make sense of mathematically.
[…]
$S=k_{B}\,\dfrac{{{\rm E}_{\scriptscriptstyle\infty}}}{E_{\lambda}},~~~~~~{\rm with}~E_{\lambda}=\dfrac{\hbar c}{\lambda},~~\lambda=2\pi{{\rm R}_{\scriptscriptstyle\infty}},$ (29)
where as above ${{\rm E}_{\scriptscriptstyle\infty}}={{\rm M}_{\scriptscriptstyle\infty}}c^{2}$. Here, the bits are quantized energies rather than hypothetical quantized patches of spacetime. Remarkably, the above formula is identical to the Gibbons-Hawking entropy, including the overall $1/4$, if one identifies $G_{N}$ as above (2):
$S=k_{B}\dfrac{{{\rm M}_{\scriptscriptstyle\infty}}c^{2}}{E_{\lambda}}=k_{B}\left({{\rm M}_{\scriptscriptstyle\infty}}c^{2}\right)\left(\dfrac{2\pi{{\rm R}_{\scriptscriptstyle\infty}}}{\hbar c}\right)=\dfrac{k_{B}}{4}\dfrac{A}{\ell_{p}^{2}},$ (30)
where in the above equation $\ell_{p}=\sqrt{\hbar G_{N}/c^{3}}$ and $G_{N}$ is given by (2). The analog of the Hawking temperature $T\equiv{{\rm T}_{\scriptscriptstyle\infty}}$ as usual follows from $d{{\rm E}_{\scriptscriptstyle\infty}}={{\rm T}_{\scriptscriptstyle\infty}}dS$ which gives the simple result
$k_{B}{{\rm T}_{\scriptscriptstyle\infty}}=E_{\lambda}.$ (31)
[…]
Let us point out that the above interpretation of the Gibbons-Hawking entropy also applies to the Bekenstein-Hawking entropy of a black hole. The reason is that (2) is the correct relation between the mass $M_{\bullet}$ of the black hole and its radius $R_{\bullet}$ if one makes the replacement ${{\rm M}_{\scriptscriptstyle\infty}},{{\rm R}_{\scriptscriptstyle\infty}}\to M_{\bullet},R_{\bullet}$:
$G_{N}=\dfrac{c^{2}}{2}\dfrac{R_{\bullet}}{M_{\bullet}}.$ (32)
From this perspective, $G_{N}$ is a universal constant which is the same for every black hole in the Universe. Repeating the above arguments for the Bekenstein-Hawking entropy,
$S_{BH}=\dfrac{k_{B}}{4}\dfrac{A}{\ell_{p}^{2}}=k_{B}\left(\dfrac{\pi c^{3}}{\hbar}\right)\dfrac{R_{\bullet}^{2}}{G_{N}}=k_{B}\,\dfrac{E_{\bullet}}{E_{\lambda}},~~~~~~{\rm with}~E_{\bullet}=M_{\bullet}c^{2}~~{\rm and}~~E_{\lambda}=\dfrac{\hbar c}{2\pi R_{\bullet}},$ (33)
where we have used (32) in the second equation. The Hawking temperature $T_{H}$ of the black hole is again given by (31), namely $k_{B}T_{H}=E_{\lambda}$.
III Induced renormalization group flow for Newton’s constant
III.1 Effective RG flow for Newton’s constant based on the RG flow for $\rho_{\rm vac}$}.
The results of the last section apply whatever $\rho_{\rm vac}$} is. In this section we use the expression (3) to obtain further results. For simplicity of subsequent discourse let us rewrite equations (19),(20) explicitly for $d=3$:
$\dfrac{\ddot{a}}{a}=\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{8\pi G_{N}}{3}{\rho_{\rm vac}}.$ (34)
In the expression (3) for $\rho_{\rm vac}$}, by dimensional analysis $\mathfrak{g}$ is a dimensionless coupling constant. We assume $m_{\rm z}$ is the physical renormalized mass and ignore its potential RG flow. There are good reasons to expect that $\mathfrak{g}$ has an RG flow, based on the comments in ALrhovac1 ; ALrhovac2 ; LeClairUniverse and equation (6) where $\alpha=e^{2}/4\pi\hbar c$ is the fine structure constant with a known RG flow coming from QED. If we associate $\mathfrak{g}$ with a marginal coupling, $\mathfrak{g}$ could be marginally relevant or irrelevant. Let us assume the following beta function, where $\mu$ is an energy scale, and increasing $\mu$ corresponds to an RG flow to higher energies [^15] The factor of $1/2\pi$ is motivated by standard definitions of fine-structure constants $\alpha$, such as $e^{2}/(4\pi\hbar c)$ for QED. If $\mathfrak{g}=\alpha=\dfrac{g^{2}}{4\pi\hbar c}$ where $g$ is a gauge coupling with beta-function $\beta_{g}=\dfrac{dg}{d\log\mu}=b\dfrac{g^{3}}{16\pi^{2}}$, then $\beta_{\mathfrak{g}}=b\mathfrak{g}^{2}/2\pi$. For pure QED with $N_{f}$ species of Dirac fermions with $g=e$, $\mathfrak{g}_{0}\approx 1/137$ and $b=4N_{f}/3$. For $\lambda\phi^{4}$ theory, if $\mathfrak{g}=\lambda$ then $b=9/8\pi$. See Peskin .
$\mu\partial_{\mu}\mathfrak{g}=\dfrac{b}{2\pi}\mathfrak{g}^{2}.$ (35)
The above beta function is typically just the 1-loop approximation, however we will argue below, based on cosmological data, that the coupling $\mathfrak{g}$ is presently small so that a 1-loop approximation is justified. We assume $\mathfrak{g}>0$ for a positive $\rho_{\rm vac}$} so that we are dealing with de Sitter space. For $\mathfrak{g}>0$, if $b$ is positive the marginal coupling $\mathfrak{g}$ increases at higher energies, namely it is marginally irrelevent, whereas if $b$ is negative $\mathfrak{g}$ flows to zero in the UV, i.e. is asymptotically free. Let $\mathfrak{g}_{0}=\mathfrak{g}(\mu_{0})$ where $\mu_{0}$ is the cosmological energy scale today at time $t_{0}$. Integrating the beta function is straightforward:
$\dfrac{\mathfrak{g}(\mu)}{\mathfrak{g}_{0}}=\dfrac{1}{1-\widehat{b}\log(\mu/\mu_{0})},~~~~~\widehat{b}\equiv\dfrac{b\,\mathfrak{g}_{0}}{2\pi}.$ (36)
The RG parameter $\widehat{b}$ is the primary new constant for applications to cosmology. As we will see, based on cosmological data, we will propose $\widehat{b}>0$, corresponding to a marginally irrelevant coupling, and unless otherwise stated, below $\widehat{b}$ is positive. The pole at $1-\widehat{b}\log(\mu/\mu_{0})=0$ is commonly referred to as a Landau pole, and signifies the theory is not UV complete. It will play an important role below where we will return to addressing its significance.
From the expression (3) one has
$\dfrac{{\rho_{\rm vac}}(\mu)}{{\rho_{\rm vac}}(\mu_{0})}=\dfrac{\mathfrak{g}_{0}}{\mathfrak{g}(\mu)},$ (37)
since we assume $m_{\rm z}$ is already renormalized to the physical mass. Thus $\rho_{\rm vac}$} in (34) should be replaced by ${\rho_{\rm vac}}(\mu)$. In order to express this equation in terms of ${\rho_{\rm vac}}={\rho_{\rm vac}}(\mu_{0})$ today, it is meaningful to incorporate the RG flow into an induced flow for Newton’s constant:
$\left(\dfrac{\dot{a}}{a}\right)^{2}={\cal H}(t;\mu)^{2}\equiv\dfrac{8\pi{\cal G}(\mu)}{3}{\rho_{\rm vac}}(\mu_{0}),$ (38)
where
${\cal G}(\mu)={\cal G}(\mu_{0})\dfrac{\mathfrak{g}_{0}}{\mathfrak{g}(\mu)}={\cal G}(\mu_{0})\left(1-\widehat{b}\log(\mu/\mu_{0})\right).$ (39)
It is meaningful to incorporate the RG flow for $\rho_{\rm vac}$} into an induced flow for $G_{N}$ since we view $G_{N}$ as being determined in the far future where vacuum energy dominates and only afterwards incorporate adding matter and energy as we will do below. If one interprets ${\cal G}(\mu)$ in (39) as an energy scale dependent Newton’s constant, then one sees that for $\widehat{b}>0$, the strength of gravity decreases as one increases the energy scale $\mu$. At extremely low energy scales $\mu$, ${\cal G}(\mu)\to\infty$. The quantity ${\cal H}(t;\mu)$ is interpreted as the solution to Einstein’s equations with a Newton’s constant ${\cal G}(\mu)$ that depends on the energy scale $\mu$. Then based on (34) one has
$\dfrac{{\cal H}(t;\mu)}{{\cal H}(t;\mu_{0})}=\sqrt{\dfrac{{\cal G}(\mu)}{{\cal G}(\mu_{0})}}=\sqrt{1-\widehat{b}\log(\mu/\mu_{0})}.$ (40)
III.2 Tying the energy scale $\mu$ to the temperature in cosmology
The natural energy scale $\mu$ in cosmology is based on its temperature $T$. In the $\Lambda{\rm CDM}$ model this temperature is tracked by the cosmic microwave background (CMB). Thus far we have only considered vacuum energy with no radiation nor matter, and will turn to the $\Lambda{\rm CDM}$ model below. As a warm up exercise, let us assume the standard relation between temperature and the scale factor: [^16] For the $\Lambda{\rm CDM}$ model, $T_{0}=2.7K$, thus $\mu_{0}=k_{B}T_{0}=2.35\times 10^{-4}\,{\rm eV}$.
$T(t)=\dfrac{T_{0}}{a(t)},$ (41)
since as we will see below the main features will persist with the addition of matter. As is also standard, we express $a(t)$ in terms of the redshift $z$, $a(t)\equiv 1/{(1+z(t))}$, such that
$\mu/\mu_{0}=T/T_{0}=1+z(t).$ (42)
We thus identify $\mu_{0}=k_{B}T_{0}$, and $z=0$ with $\mu_{0}$. If we identify $G_{N}={\cal G}(\mu_{0})$, then
${\cal G}(\mu)=G_{N}\left(1-\widehat{b}\log(1+z)\right)=G_{N}\left(1+\widehat{b}\log(a)\right).$ (43)
Note that $G_{N}=0$ when $a=a_{\rm min}=e^{-1/\widehat{b}}$. For smaller $a(t)<a_{\rm min}$, in principle ${\cal G}(\mu)$ could change sign, which would signify a transition to anti-de Sitter space, however as we will see, $a(t)$ never drops below $a_{\rm min}$ and the logarithmic branch cut from $a=0$ is never reached.
Turning to (38), $\dot{a}/a=H$ is no longer constant since $\mu$ is time dependent according to (42). Taking the square-root we wish to solve
$\left(\dfrac{\dot{a}}{a}\right)=\pm\sqrt{1+\widehat{b}\log a(t)}~\left(\dfrac{8\pi G_{N}\rho_{{\rm vac};0}}{3}\right)^{1/2}$ (44)
where $\rho_{{\rm vac};0}$ is $\rho_{\rm vac}$} at the scale $\mu_{0}$. This additional kind of time dependence does not originate directly from Einstein’s equations but rather from the energy scale $\mu$ dependence of Newton’s constant and $\mu$ being tied to the temperature through equation (42). This leads to a significant and interesting modification of the de Sitter space solution to $a(t)$. Let us first chose the positive sign ++ on the RHS of (44) such that $\dot{a}>0$ in the future. As we will show, choosing the minus sign will be relevant for the far past. The equation (38) can be explicitly integrated, yielding a “inverted gaussian” correction to de Sitter space, which we will simply refer to as gaussian de Sitter space:
$a(t)=e^{-1/\widehat{b}}\,\exp\left[\dfrac{1}{\widehat{b}}\left(\dfrac{\widehat{b}H_{0}}{2}(t-t_{0})+1\right)^{2}\right]=\exp\left[H_{0}(t-t_{0})+\widehat{b}H_{0}^{2}(t-t_{0})^{2}/4\right].$ (45)
Above, $H_{0}$ equals $\dot{a}/a$ at the present time $t_{0}$, since the constant of integration is chosen such that $a(t_{0})=1$. One sees that near the present time, namely $t-t_{0}$ small, the second term in the exponential $\propto(t-t_{0})^{2}$ is suppressed, and if $\widehat{b}=0$ one recovers the usual de Sitter solution $a(t)=e^{H(t-t_{0})}$ with $H=H_{0}$.
The most interesting feature of the gaussian de Sitter solution (45) is that there is no singularity $a=0$ for all times. In spite of the Landau pole, the solution is still regular, since the argument of the square root, $1+\widehat{b}\log a(t)$, is always positive since $a(t)>a_{\rm min}$ for all times. This minimum value of $a(t)$ is
$a(t)>a_{\rm min}\equiv e^{-1/\widehat{b}}~~\forall t~~~~~\Longrightarrow~~z_{\rm max}=\dfrac{1}{a_{\rm min}}-1=e^{1/\widehat{b}}-1.$ (46)
The time $t_{\rm min}$ where $a(t_{\rm min})=a_{\rm min}$ is equal to
$t_{\rm min}=t_{0}-\dfrac{2}{H_{0}\widehat{b}}$. (47)
Note that as $\widehat{b}\to 0$, $t_{\rm min}\to$-$\infty$ as expected for de Sitter space. In the next section we will approximate $\widehat{b}\approx 0.02$ based on the so-call Hubble tension, which leads to a very large value for $z_{\rm max}$.
Let us now turn to the far past $t\to$-$\infty$. In ordinary de Sitter space, $\lim_{t\to-\infty}a(t)=0$, although it is known this is not a true geometric coordinate singularity. For our gaussian de Sitter space, the solution (45) is formally still valid for $\infty<t<t_{\rm min}$, and thus valid for all times. In fact
$\lim_{t\to\pm\infty}a(t)=\infty.$ (48)
This is a consequence of a stronger relation between the past and future. One can see from the solution (45) that it has the symmetry
$a(t)=a(-t+2t_{\rm min}).$ (49)
It is important to note that the above symmetry is independent of the present time $t_{0}$. At the time $t_{\rm min}$,   $a(t_{\rm min})=a_{\rm min}$ is self-dual. Recall that at the time $t_{\rm min}$ with scale factor $a_{\rm min}$, the effective Newton constant ${\cal G}(\mu)$ vanishes such that the solutions of (44) with ++ verses - agree, thus for $t<t_{\rm min}$ one should take the solution with the minus sign, where $a(t)$ decreases.
To summarize, the solution (45) is valid for all times $\infty<t<\infty$ where $a(t)>a_{\rm min}$. The complete history of such a Universe is that in the far past $t\to$-$\infty$ the scale factor $a(t)$ is infinitely large and starts to compress down to $a_{\rm min}$ which occurs at a time $t_{\rm min}$. Beyond this time, $t>t_{\rm min}$, the Universe expands until the scale factor $a(t)$ is again infinite at $t=+\infty$. This takes an infinite amount of time, and avoids any geometric singularities at $a=0$. There is no “Big Bang” that occurs at time $t_{\rm min}$, rather at this time the Universe smoothly transitions from a compression to an expansion, and $t_{\rm min}$ is just the time when the Universe is hottest. The Big Bang is now better described as a Swing. An picturesque analogy is the harmonic pendulum with the bottom of the swing corresponding to $a_{\rm min}$, and at the present time the Universe is on the up swing. Based on (49), a consistent possibility is that this process then repeats itself, such that Universe is in a sense oscillating about its vacuum like a pendulum harmonic oscillator. At the top of the swing the Universe is at its largest. If this is the case, then the age of this gaussian de Sitter Universe is even more infinite.
IV Adding matter: the $\Lambda{\rm CDM}$ model re-examined and the Hubble Tension
Thus far we have only considered vacuum energy density $\rho_{\rm vac}$}. We gave a heuristic derivation of Einstein’s equations in this context from the gravitational Casimir effect we formulated above. The Universe also contains radiation and matter, which we view as excitations above the vacuum, which must be included in $T_{\mu\nu}$. Local conservation requires vanishing of the covariant divergence $\nabla^{\mu}T_{\mu\nu}=0$, which is automatic due to the Bianchi identities. Thus, for instance, Newton’s universal law for the gravitational force between two masses is subsumed as a consequence. At this point one must deal with the usual Einstein’s equations except with an energy dependent Newton’s constant ${\cal G}(\mu)$:
$\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{8\pi}{3}{\cal G}(\mu)\,\rho_{\rm total},$ (50)
with ${\cal G}(\mu)$ defined in (43). For the moment, let us not tie the energy scale to $\mu$ to $z(t)$ as we did in (42). It is conventional to express (50) as follows
$H^{2}\equiv\left(\dfrac{\dot{a}}{a}\right)^{2}=H_{0}^{2}\left(1+\widehat{b}\log a\right)\left(\dfrac{\Omega_{\rm rad}}{a^{4}}+\dfrac{\Omega_{\rm m}}{a^{3}}+\Omega_{\Lambda}\right),$ (51)
where by definition $a(t_{0})=1$ at the present time $t_{0}$, $H(t_{0})=H_{0}$, and $\Omega_{\Lambda}$ is the $\rho_{\rm vac}$} contribution. The zero curvature ($k=0)$ motivated in Section II implies $\Omega_{\rm rad}+\Omega_{\rm m}+\Omega_{\Lambda}=1$. In the $\Lambda{\rm CDM}$ model one ignores radiation, i.e. $\Omega_{\rm rad}\approx 0$, as this is a well-justified approximation during this epoch.
There is a deepening discrepancy in the $\Lambda{\rm CDM}$ model based on relatively recent astrophysical measurements referred to as the “Hubble tension”. The ideas in this article offer a potential resolution which we now present. As usual, let $H_{0}=\dot{a}/a$ be the Hubble constant today. It should be a fixed constant regardless of the manner in which it is measured. The discrepancy arises from two very different kinds of measurements. The first comes from “local” measurements based on Type Ia supernovae, where typical redshifts are relatively low, in the range z=0.020.15z=0.02–0.15 (SHOES SHOES .). The other determination of $H_{0}$ is based on Cosmic Microwave Background data at much higher redshift $z=1100$. (Planck Planck ). The two contradictory values reported are
$H_{0;{\rm CMB}}=67.4\pm 0.5~{\rm km/s/Mpc},~~~~~H_{0;{\rm SN}}=73.0\pm 1.0~{\rm km/s/Mpc},$ (52)
which differ significantly enough for some prominent researchers to question the $\Lambda{\rm CDM}$ model, with suggestions that this could signify beyond standard model physics HubbleTensionRef . In the latter it was suggested that this could be due to a variable Dark Energy component, but no specific theoretical model was proposed.
Our interpretation of the discrepancy in these two values for $H_{0}$ in (52) is the following. Consider an observer making measurements of $H(t)$ at an earlier time, where their measurements are in a small range of higher redshift $z$ than today where today $z=0$. At this epoch, ${\cal G}(\mu)$ is effectively lower according to (43). If this observer fits their data to (51) they will predict a lower value of $H_{0}$ compared to a fit based on data taken at $z\approx 0$. The CMB measurements probe the state of the Universe at this earlier time, thus though CMB measurements are made at the present time $t_{0}$, they reflect a hypothetical observer making measurements at an earlier epoch where $z\approx 1100$. We thus propose based on (40)
$\dfrac{H_{0;{\rm CMB}}}{H_{0;{\rm SN}}}\approx\sqrt{1-\widehat{b}\log(1+z)},~~~~{\rm for}~z=z_{\rm CMB}=1100,$ (53)
where we have set $z\approx 0$ for the supernovae measurements. Based on the measured values in (52), the above leads to $\widehat{b}\approx 0.02$.
Let now turn to explicitly solving (51) in “real time”, namely with ${\cal G}(\mu(t))$ given in (42), as we did for pure vacuum energy with the result (45). Before even integrating the equation (50), one can immediately see that there are no real solutions unless $1+\widehat{b}\log a>0$. Thus there is still a minimal scale factor $a(t)>a_{\rm min}$ which is the same as in (46) and attributed to the Landau pole. This is a robust conclusion, regardless of whether $\rho_{\rm rad}$ is incorporated as the dominant form of energy in the very early Universe. The above value for $\widehat{b}$, leads to a very high value for the maximal red shift: [^17] This is not an unreasonable value for $\widehat{b}$, in that $\mathfrak{g}_{0}=2\pi\widehat{b}/b=0.12/b$, and typically $b={\cal O}(1)$. For comparison, for QED $b=4N_{f}/3$ for $N_{f}$ species of Dirac fermions, and if $\mathfrak{g}_{0}=1/137$ then $\widehat{b}=.002N_{f}$ which is about 10 times smaller than in (54) and $z_{\rm max}\approx 10^{500}$ is far beyond the Planck scale.
$\widehat{b}\approx 0.02~~~\Longrightarrow~~a_{\rm min}\approx 2\times 10^{-22},~~~~z_{\rm max}\approx 5\times 10^{21}.$ (54)
This is far above what has been directly probed by the CMB, the latter being $z_{\rm CMB}=1100$. Thus our proposed RG flow for Newton’s constant implies only small logarithmic corrections to the $\Lambda{\rm CDM}$ model. For instance it doesn’t significantly alter what was previously considered as the age of the Universe, as we will show below.
To further justify the last statement, let us turn to the $\Lambda{\rm CDM}$ model where radiation $\Omega_{\rm rad}=0$ as a good approximation,. In order to probe potential singularities at early times where the matter dominates due to the $1/a^{3}$ in (50), let us neglect the $\Omega_{\Lambda}$ term as an approximation. Then one wishes to integrate the equation
$\left(\dfrac{\dot{a}}{a}\right)^{2}=H_{0}^{2}\,\left(1+\widehat{b}\log a\right)\,\dfrac{\Omega_{m}}{a^{3}}.$ (55)
When $\widehat{b}=0$, this is easily integrated:
$a(t)^{3/2}=\tfrac{3}{2}H_{0}\sqrt{\Omega_{m}}\left(t-t_{0}+\dfrac{2}{3H_{0}\sqrt{\Omega_{m}}}\right),$ (56)
where $a(t_{0})=1$ is an initial condition. There exists a time $t_{\rm min}$ where aa=0: [^18] In the cosmology literature the convention is to shift $t\to t+t_{\rm min}$ such that $a=0$ at the shifted time $t=0$. The difference $t_{0}-t_{\rm min}$ essentially represents the age of the Universe if one assume the Universe came into existence at time $t_{\rm min}$, which is questioned in this article.
$a(t_{\rm min})=0~~{\rm for}~~t_{\rm min}=t_{0}-\dfrac{2}{3H_{0}\sqrt{\Omega_{m}}}.$ (57)
For $\widehat{b}\neq 0$ the equation (55) can also be explicitly integrated. As for the pure vacuum energy case of the last section, taking the square-root of (55) introduces a $\pm$ sign as in (44), where +/+/- corresponds to the far future/past. For the ++ sign the solution $a(t)$ can be expressed implicitly in terms of the imaginary error function ${\rm erfi}(x)\equiv{\rm erf}(ix)/i$:
${\rm erfi}\sqrt{\dfrac{3}{2\widehat{b}}\left(1+\widehat{b}\log a(t)\right)}=(t-t_{0})H_{0}\,\,\sqrt{\dfrac{3\widehat{b}\,\Omega_{m}}{2\pi}}\,e^{3/2\widehat{b}}+{\rm erfi}\sqrt{\dfrac{3}{2\widehat{b}}}$, (58)
where again we have imposed $a(t_{0})=1$. The argument of the $\rm erfi$} function on the LHS must be real otherwise the $\rm erfi$} function is imaginary. Thus due to the square-root branch cut $\sqrt{1+\widehat{b}\log a}$ inside $\rm erfi$}, there is a minimal value of $a(t)$ again given by (46). This minimal value of $a(t)$ occurs at a time $t_{\rm min}$ where $a(t_{\rm min})=a_{\rm min}$, which is determined by when the LHS of the above equation is zero since ${\rm erfi}(0)=0$:
$a(t_{\rm min})=a_{\rm min}=e^{-1/\widehat{b}}$ (59)
where
$t_{0}-t_{\rm min}=\dfrac{1}{H_{0}}e^{-3/2\widehat{b}}\sqrt{\dfrac{2\pi}{3\widehat{b}\,\Omega_{m}}}{\rm erfi}\sqrt{\dfrac{3}{2\widehat{b}}}=\dfrac{2}{3H_{0}\sqrt{\Omega_{m}}}\left(1+\tfrac{1}{3}\,\widehat{b}+\tfrac{1}{3}\,\widehat{b}^{2}+\tfrac{5}{9}\,\widehat{b}^{3}+\ldots\right).$ (60)
The solution (58) is again valid for all times $\infty<t<\infty$ due to a symmetry between the far past and far future. Using ${\rm erfi}(-x)=-{\rm erfi}(x)$, one can show that equation (49), which equates $a(t)$ with $a(-t+2t_{\rm min})$, remains valid with $t_{\rm min}$ given in (60), as does (48).
Based on the above analysis of our scenario, let us comment on the possible implications for the very early Universe. Henceforth, by “early Universe” we mean around the time $t_{\rm min}$, where the expansion phase begins. Recall in our model, time $t$ actually extends to $t=$-$\infty$, and this is the earliest Universe. The salient features of the evolution of $a(t)$ from $t=$-$\infty$ to $t=+\infty$ is essentially the same as for the pure vacuum energy case, described in the last paragraph of the last section. Namely at $t=$-$\infty$, $a(t)=\infty$ and starts to decrease until it reaches $a_{\rm min}$ at time $t_{\rm min}$, then begins to increase back to $\infty$ as $t\to+\infty$. At no time in this evolution is $a=0$, again avoiding geometric curvature singularities. There is no singular Big Bang at any time, rather at time $t_{\rm min}$ the universe is hottest since the redshift $z$ is at its maximum. The quantity $t_{0}-t_{\rm min}$ does not represent the age of the Universe, but rather the time elapsed since the scale factor was at its minimum. The age of the Universe is actually infinite, which we find appealing, since the Universe does not originate from an incomprehensible singularity at $a=0$. As described above for the case of only Vacuum energy, the time evolution of the Universe is better described as a back and forth swing of a pendulum, where the bottom of the swing occurs at $a(t_{\rm min})=a_{\rm min}$, and we are currently on an upswing.
The details of this evolution are very sensitive to the value of $\widehat{b}$ due to the exponentials in (54). For purposes of illustration, let us assume $\widehat{b}=0.02$ as inferred from the Hubble tension. First, one sees from (60) that for small $\widehat{b}$ this amounts to a small corrections to $t_{0}-t_{\rm min}$ which is normally associated to the age of the Universe. If our estimate of $\widehat{b}$ in (54) is approximately correct, then this value for $z_{\rm max}$ is deeply into the radiation dominated era, since the value of $z=z_{\rm eq}$ where matter and radiation are equivalent in density is only about $z_{\rm eq}\approx 3400$. At times much earlier than during the $\Lambda{\rm CDM}$ era, equation (42) is not necessarily valid, however if we simply assume it extends to the Planck scale with $T_{0}=2.7$K, then $z_{\rm planck}\approx 5\times 10^{31}$, which is much higher than $z_{\rm max}$ by 10 orders of magnitude. This indicates that the Planck scale plays no role in determining $z_{\rm max}$, which is to say quantum gravity does not play a role.
For our perspective on gravity, earlier times $t<t_{\rm min}$ where $a(t)<a_{\rm min}$ do not exist in our model. In the standard cosmology this earlier era is thought to be described by inflation. Inflation is postulated as a way to get from the singularity at $a=0$ to a hot Big Bang, however it is not yet universally accepted as being a complete theory. Very large numbers are naturally involved since the singularity at $a=0$ formally corresponds to $z=\infty$. Typical models require inflation to last for 50-60 e-foldings to inflate the Universe to the proper initial conditions from $a\approx 0$, then transition to a re-heating scale to bring up the temperature to that of a hot Big Bang, and this rapid expansion makes the Universe flat and homogeneous. The range of redshift $z$ over which inflation hypothetically occurs is strongly model dependent, especially on what the re-heating temperature is, and the duration of inflation, marking its beginning and end, can range from $10^{22}<z_{\rm inflation}<10^{53}$, where the upper limit can even be significantly larger than $z_{\rm Planck}$. [^19] During hypothetical inflation, the redshift $z$ is not given by the formula (42), but based on the definition of the scale factor $a=1/(1+z)$. It is noteworthy that at the lower end of this range based on lower reheating temperatures, namely where inflation ends at $z_{\rm inflation}=10^{22}$, is close to our $z_{\rm max}=5\times 10^{21}$. Since this inflationary epoch does not actually exist in our model, one should question whether inflation is necessary in the standard model of cosmology. At $t=t_{\rm min}$ the Universe is already expanded since $a(t)>a_{\rm min}$, and the Universe is already flat with $k=0$. In Section II we explained why the spatial curvature $k$ is naturally zero since we view Gravity as arising from vacuum energy in flat Minkowski space. The value of $a_{\rm min}=e^{-1/\widehat{b}}$ perhaps plays a role equivalent to hypothetical inflation models since it represents roughly 50 e-foldings since $e^{1/\widehat{b}}\approx e^{50}$ for our estimate of $\widehat{b}\approx 0.02$.
V Closing Remarks
Having already summarized the unconventional perspective on Gravity developed in this article and its implications in the Abstract and Introduction, we close with some suggestions for further investigations and discuss some open questions. There are currently many open avenues for testing (and falsifying) our main proposals.
$\bullet$   Above we considered implications for the cosmology based on the $\Lambda{\rm CDM}$ model. The energy scales involved are relatively low since $\mu_{0}=k_{B}T_{0}=2.35\times 10^{-4}\,{\rm eV}$ for $T_{0}=2.7$ K. If our proposed energy dependent Newton’s constant ${\cal G}(\mu)$ in equation (43) is correct, then it is in principle possible to confirm it in some bench-top Cavendish like experiments as a function of temperature. Suppose such experiments are carried out in a small range of temperature $\Delta T$ about a fixed temperature $\widetilde{T}$. From (39)
${\cal G}(T)={\cal G}(\widetilde{T})\left(1-\widehat{b}\log(T/\widetilde{T})\right)\approx{\cal G}(\widetilde{T})\left(1-\dfrac{\widehat{b}}{\widetilde{T}}\,\Delta T\right).$ (61)
As discussed above, the effective Newton’s constant decreases with increased temperature. If $\widetilde{T}=300$ K is room temperature, then $\widehat{b}/\widetilde{T}\approx 6\times 10^{-5}$ per degree change $\Delta T$ if $\widehat{b}=0.02$, where this value of $\widehat{b}$ was inferred from the Hubble tension in (53). There are relatively few experiments along these lines, going back as far as 1923 Shaw ; Dmitriev ; Liangzao , wherein weight is measured as a function of temperature for various metals around room temperature. These experiments indeed consistently measure a decrease in the gravitational force as the temperature is increased, as our model predicts. Based on the ideas in this article, this decrease should be independent of the material, in this case a metal, since the Equivalence Principle is built into the formalism, whereas experiments find slightly different relative changes to $G_{N}$ depending on the metal. This could be due to experimental uncertainties as a result of buoyancy, thermal expansion, and other factors which are difficult to take into account. In any case a linear fit of this experimental data in $\Delta T$ according to (61) yield variations of about $10^{-5}$ per degree, which is roughly consistent with our estimate of $\widehat{b}/\widetilde{T}$ for $\widehat{b}\approx 0.02$ and $\widetilde{T}=300$K. [^20] The article Liangzao does not attempt a fit to their data. We made a crude fit of their tabular results to the functional form (61) and found $\widehat{b}/\widetilde{T}$ to be $0.8\times 10^{-5}$ for gold and $1.4\times 10^{-5}$ for nickel, however a closer inspection and interpretation of this data beyond our crude fit is warranted. At lower reference temperatures $\widetilde{T}$, the effect is larger and this motivates carrying out such Cavendish type experiments at lower temperatures.
$\bullet$   The induced RG flow for Newton’s constant that we proposed as a consequence of the RG flow of the coupling $\mathfrak{g}$ in $\rho_{\rm vac}$} in (3) led to an avoidance of a geometric curvature singularity at a scale factor $a=0$. There are various theorems asserting that such singularities are unavoidable in cosmology and black holes Penrose ; HawkingPenrose . The latter theorems assume a constant Newton’s constant $G_{N}$, and should be revisited to account for an effective energy scale dependent $G_{N}$ as proposed in this article. This could potentially lead to a resolution of the singularity at the origin of the Schwarzschild black hole solution, in a way analogous to the minimal scale factor $a_{\rm min}$ above.
$\bullet$   If the varying $G_{N}$ proposed in this article is correct, it is testable by other kinds of measurements. Let us speculate on some possible directions or topics for further study. We already mentioned above that in principle this varying $G_{N}$ can be measured in bench-top Cavendish experiments as a function of temperature, and that the very few existing experiments carried out at room temperature are roughly consistent with the parameters we inferred from the Hubble tension. For cosmology, our formula (39) for the scale dependent Newton’s constant ${\cal G}(\mu)$ implies that the effective Newton’s constant is effectively time dependent if one ties $\mu$ to the time-dependent scale factor as in (42). There are known experimental bounds on the time variation of $G_{N}$, usually reported as $\dot{G}_{N}/G_{N}$. Based on (43) one has
$\dfrac{\dot{G}_{N}}{G_{N}}=H\,\widehat{b}$, (62)
where as above $H=\dot{a}/a$. The strongest experimental constraints come from pulsar timing Taylor which constrains $\dot{G}_{N}/G_{N}<2\times 10^{-12}/{\rm yr}$. LIGO provides weaker constraints $10^{-9}/{\rm yr}$ LIGO . These pulsars exist at relatively low redshift $z$. Thus taking $H=H_{0}$ in the above equation (62), one finds $\dot{G}_{N}/G_{N}\approx 7\times 10^{-11}\times\widehat{b}\,/{\rm yr}$. For $\widehat{b}\approx 0.02$, as we proposed based on Hubble tension data, $\dot{G}_{N}/G_{N}\approx 1.4\times 10^{-12}\,/{\rm yr}$ which is very close to the bound based on pulsars. There are many other testable predictions in cosmology in addition to the Hubble tension and those mentioned above. A varying $G_{N}$ may shift CMB acoustic peaks due to modified expansion at $z\sim 1100$, which is testable with Planck or the Simons Observatory data. The minimal scale factor $a_{\rm min}\sim 10^{-22}$ in eq. (46) could produce a primordial gravitational wave spectrum detectable by LIGO or eventually the space based interferometer LISA. A weaker $G_{N}$ at higher redshift $z$ could delay stellar collapse, favoring primordial black holes. In fact, based on (32), for a weaker $G_{N}$, black holes are smaller for a given mass. Very recently, supermassive black holes, with mass on the order of 40 million solar masses have been discovered using the James Webb Space Telescope JWST which were formed at an earlier time than expected, and this may be an interesting topic for further study in the context of a varying $G_{N}$.
$\bullet$   As stated in the Introduction, although the formula (3) is well-motivated by the work ALrhovac1 ; ALrhovac2 , the QFT underlying this formula for $\rho_{\rm vac}$} has not been precisely identified, in particular the zeron with mass $m_{\rm z}$ has not yet been identified as a known particle, though Majorana fermions are viable candidates. This underlying QFT is incomplete in the UV since it relies on the RG beta function (35) for a marginally irrelevant coupling, where the incompleteness is manifested as a Landau pole at very high energies. However we showed that the time evolution of the scale factor $a(t)$ smoothly traverses this pole through a swing rather than a bang. The UV completeness issue here is certainly more tractable than the UV incompleteness of quantum gravity, since it is a UV issue of QFT in flat Minkowski space where such problems are better understood. Landau poles are commonplace in elementary particle physics and condensed matter physics, since they are generic to marginally irrelevant perturbations. See for instance LandauPole1 ; LandauPole2 ; LandauPole3 and references therein. Although QED itself is not relevant to this article, let us mention that it also is not UV complete due to a Landau pole, however in this context it occurs at much higher energies than our proposed $z_{\rm max}$, and is thus commonly viewed as a purely academic issue, and is furthermore complicated by chiral symmetry breaking. (See footnote 12).

### v3

@@
 Gravity remains the least understood of the fundamental forces, in spite of the discoveries of General Relativity (GR) and Quantum Mechanics (QM) over 100 years ago, this year being recognized as the 100-th anniversary of a complete theory of QM due to Schrödinger and Heisenberg, although the birth of QM goes back to Planck’s black body studies at finite temperature.
-Classical singularities at the center of black holes and at the time of the Big Bang have not yet been understood and resolved in classical GR.
-Many physicists expect that such singularities will be resolved by a theory of quantum gravity, but it’s fair to say that a complete theory of quantum gravity has not been forthcoming, although string theory offers some promise.
-For these reasons it is worthwhile to pursue new ideas on the origin or emergence of Gravity itself, especially if such ideas suggest that perhaps Gravity doesn’t need to be quantized in the first place.
+Classical singularities at the center of black holes and at the time of the so-called Big Bang have not yet been understood and resolved in classical GR.
+Many physicists expect that such singularities will eventually be resolved by a theory of quantum gravity, but it’s fair to say that a complete theory of quantum gravity that can address these fundamental issues has not been forthcoming.
+String theory still offers some promise and is the most well-motivated since it provides a consistent scattering theory of gravitons and more.
+Nevertheless, for these reasons it is worthwhile to pursue new ideas on the origin or emergence of Gravity itself, especially if such ideas hint that perhaps Gravity doesn’t need to be quantized in the first place, or at the very least is not yet needed to explain currently observable phenomena.
 Such a theory should be able to explain the value of Newton’s constant $G_{N}$ itself.
-In this article we explore the question “Can Quantum vacuum energy be the origin of Gravity?” [^2] This article was initially intended simply as a theoretical exercise in the form of an essay to address this question, but we found it can lead to a fruitful new perspective on Gravity more generally, and it is not without some predictions, if the basic ideas continue to hold.
-This perspective is based on the idea that perhaps we can learn more about the foundations of Gravity by considering the cosmological far future, in contrast to the far past where much less is understood, in particular the origin of the Big Bang and possibly inflation.
+Like Gravity, the Vacuum $|0\rangle$ is omnipresent.
+The Vacuum “knows” the physical properties, such as mass, quantum numbers, and interactions of any particle excitations created above it, and is perhaps the only entity that exists everywhere in the Universe and for all times.
+The existence of the Quantum Vacuum guarantees that the allowed particle excitations above it, such as electron/positron pairs, are identical throughout the Universe, because they already exist virtually in both a physical and mathematical sense.
+Everything else we observe are quantum excited states of the Vacuum above the ground state it represents, including people, and are typically unstable.
+Nature does not abhor a Vacuum as Aristotle first stated, rather it depends on it.
+[^2] To paraphrase with translation a recognizable philosopher, If you look deeply enough into the Void, It eventually looks back at you.
+So it is reasonable to suppose that the Vacuum, in particular its ground state energy density $\rho_{\rm vac}$}, is at the origin of Gravity.
+If the only way to measure the ground state energy $E_{0}$ is with Gravity, then reversing the argument, perhaps quantum vacuum energy is the origin of Gravity.
+[^3] In a certain sense, the old aether is back as the Vacuum itself.
+Deceptively simple as it may appear, to study quantum field theory (QFT) at finite temperature $T$ with periodic spatial boundary conditions, one necessarily needs to study quantum fields in curved euclidean space where the euclidean time is compactified to a circle with circumference $\hbar c/k_{B}T$.
+In one spatial dimension the euclidean spacetime is a torus, and the problem of $d=1$ conformal QFT on a torus is completely solved, even though it is a non-trivial problem involving modular invariance, etc.
+From this point of view, temperature is a fundamental link between geometry and QFT.
+For the same reason that one does not necessarily have to quantize the heat bath to study finite temperature QFT, perhaps one does not have to quantize Gravity.
+If the Vacuum and its energy density is the origin of Gravity itself, then since the Vacuum is intrinsically Quantum, one ponders whether quantizing gravity is essentially redundant.
+Henceforth, we set aside the deep question as to whether Gravity ultimately needs to be quantized since, as we will see, a resolution of the singularity at the origin of the so-called Big Bang can be understood without it.
+[^4] We take “quantum gravity” to mean quantization of the metric gravitational field itself, which necessarily implies spin-2 massless gravitons.
+Other arguments that attempt to prove that gravity must be quantized are not based on gravitons, but rather on the principles of superposition and unitarity in QM.
+These arguments attempt to prove that a purely classical gravitational field leads to inconsistencies with the rest of the quantized world.
+In a somewhat fanciful thought experiment put forward by Feynman Feynman , imagine a massive ball that becomes entangled with an electron, where the electron is in a superposition of two states with different paths, thus the ball becomes a superposition of states with different locations.
+This would imply the gravitational field of the ball is in a superposition, and this could in principle be detected by interacting gravitationally with a second ball.
+These arguments rely on unitarity of QM, and whether the Universe is an open or closed system is debatable.
+For instance, in the search for a de Sitter/CFT correspondence, it has been proposed that the CFT should be non-unitary Strominger .
+Recently a potential loop-hole in Feynman’s thought experiment has been presented Oppenheim .
+This does not preclude the existence of classical gravitational waves that have been measured, as they can be viewed as classical perturbations of the Vacuum as an elastic medium.
+For purposes of illustration, consider a universe which only consists of a single harmonic oscillator in its ground state with energy $E_{0}=\hbar\omega_{0}/2$, in addition to an imagined observer.
+Since classical mechanics is unaffected by shifts of the potential energy by an arbitrary constant, $E_{0}$ is not measurable in classical physics.
+If all that exists is the harmonic oscillator in empty space, then $E_{0}$ also cannot be measured in a single measurement based on QM, but can in principle be measured if coupled for instance to photons with repeated measurements.
+There are two known general ways to measure $E_{0}$.
+First, one can couple the harmonic oscillator to a generic heat bath at temperature $T$.
+Then one can in principle infer $E_{0}$ from the high temperature behavior of the partition function or its logarithm which is the free energy.
+It’s important that this is based on the free energy and not its derivatives, and thus it is not invariant under shifts of the potential energy, and this avoids some fine-tuning issues.
+The second manner to measure $E_{0}$ is to couple the oscillator to Gravity, since gravity originates from all the matter-energy in the Universe and is sensitive to its ground state energy.
+The fact that the ground state energy can be measured either by coupling to a heat bath or gravity suggests parallels between thermodynamics and gravity that have already been recognized, in particular for the entropy of black holes.
+If our imagined observer of the single harmonic oscillator thinks about natural units, they would be based on $\hbar$, $c$, and $E_{0}$ which are complete, as they lead to a fundamental length, time and mass.
+Based on the above remarks, the initial perspective on Gravity taken in this article is based on the idea that perhaps we can learn more about the foundations of Gravity by considering the cosmological far future, in contrast to the far past where much less is understood, in particular the origin of the Big Bang and possibly inflation.
 The late Universe is dominated by so-called Dark Energy, thus one should first attempt to understand the full implications of this in a Universe where Dark Energy is its sole component.
+Below we will refer to such a hypothetical universe as the dark-universe, and it should be viewed as a skeleton of a model of the Universe where there are no excitations over its ground state.
 By assumption, we will equate Dark Energy with the quantum vacuum energy density $\rho_{\rm vac}$} computed in quantum field theory (QFT) in flat local Minkowski space, defined as
@@
 The gravitational implications of the quantum field theoretic properties of $\rho_{\rm vac}$} will be then explored.
+Thus this article essentially explores the question “Can Quantum vacuum energy be the origin of Gravity?” The ideas presented here are most closely related to Andrei Sakharov’s idea on “induced gravity” Sakharov ; Visser , however they differ in many important details.
+[^5] The parallels with Sakharov’s induced gravity ideas were only pointed out to us after the first version of this article was made available on arXiv.
+Sakharov’s original article is extremely short with very little equations and is difficult to penetrate.
+A nice explanation can be found in MisnerThornWheeler page 426, where it is explained how Newton’s constant $G_{N}$ is viewed as a kind of elastic modulus.
+For instance our model does not require a UV cut-off in momentum space as Sakharov’s theory does, since this UV divergence is controlled in the QFT through renormalization, and this renders $\rho_{\rm vac}$} well-defined both physically and mathematically.
+This article can thus be viewed as applying modern ideas of the renormalization group to Sakharov’s idea to deal with the cut-off, which were not formulated until after Sakhavov’s work by Wilson KWilson , and the resulting renormalization group flow for $\rho_{\rm vac}$}.
+The theoretical framework is still semi-classical gravity where $T_{\mu\nu}$ for Dark Energy is identified with its vacuum expectation value.
+We are unaware of applications of Sakharov’s ideas to the Hubble tension in the literature, nor its prediction of the minimal scale factor $a_{\rm min}$ described below.
+The main assumptions in this article are rather conservative and minimal, and it’s useful to itemize them here in order to understand how they fit into the existing literature.
+Assumption 1.
+Dark Energy, or equivalently the Cosmological Constant, is equated with quantum vacuum energy density in flat Minkowski space, namely the vacuum expectation value of the energy-momentum tensor (1).
+This quantity is finite and well-defined in any QFT and studied in some detail recently for 4 dimensional spacetime ALrhovac1 ; ALrhovac2 .
+There we obtained the formula
+${\rho_{\rm vac}}=\dfrac{3}{4}\dfrac{c^{5}}{\hbar^{3}}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}}$ (2)
+where $m_{\rm z}$ is the physical (renormalized) mass of the lightest massive particle, and $\mathfrak{g}$ is a dimensionless interaction coupling constant, or a function thereof if higher order corrections in perturbation theory are included.
+The above quantity refers to the vacuum energy density for all known particles, and the fact that the lightest mass particle $m_{\rm z}$ appears is based on bootstrap principles.
+This thus assumes there is a single particle mass scale that determines all other particle masses.
+This is actually the situation in the Standard Model of particle physics where all masses are thought to arise from the single scale through the Higgs mechanism.
+This assumption is even more correct if there is a Grand Unified Theory of particle physics.
+If the Vacuum and its energy density $\rho_{\rm vac}$} were based on two or more completely independent and decoupled QFT’s with different fundamental mass scales, then the formula (2) doesn’t obviously apply since it becomes undetermined what $m_{\rm z}$ actually is.
+Thus we have to assume all particles can interact with each other, otherwise one is dealing with separate, distinct universes.
+The formula (2) is not ad hoc, and was well-motivated in ALrhovac1 ; ALrhovac2 .
+Since this formula is central to our study, its underpinnings are reviewed and explained below in Section IVA.
+It is not necessary for our purposes at this preliminary stage to definitively identify the $m_{\rm z}$-particle as a known particle, however we can constrain its mass, equation (5), and it’s coupling constants based on Hubble tension data, and we will have more to say about its properties below.
+Assumption 2.    The coupling $\mathfrak{g}$ is a marginally irrelevant coupling with 1-loop renormalization group (RG) beta function:
+$\mu\partial_{\mu}\mathfrak{g}=\dfrac{b}{2\pi}\mathfrak{g}^{2}$ (3)
+for some positive constant coefficient $b$, where increasing $\mu$ corresponds to a flow to higher energies.
+Integrating the RG flow leads to the parameter $\widehat{b}=b\,\mathfrak{g}_{0}/2\pi$ where $\mathfrak{g}_{0}$ is the value of $\mathfrak{g}$ at the relevant energy scale today.
+The parameter $\widehat{b}$ is what modifies the Friedmann equations in the way presented below.
+Based on the Hubble tension we will constrain $\widehat{b}\approx 0.02$ which is quite small, and justifies a 1-loop approximation.
+Assumption 3.    There exists an energy scale of the Universe corresponding to a time dependent temperature $T(t)$, where as usual
+$T(t)=\dfrac{T_{0}}{a(t)}=T_{0}\,(1+z(t))$ (4)
+where $a(t)$ is the scale factor and $z$ the redshift, and $T_{0}$ is the temperature today.
+As we will see, de Sitter space has a natural temperature (29).
+We assume $T_{0}$ is the temperature of the Cosmic Microwave Background (CMB) today, $T_{0}\approx 2.7$K, however some predictions are not so sensitive to the exact value of $T_{0}$ and it could in principle be the temperature based on neutrinos which is not so different.
+We mainly focus on the $\Lambda{\rm CDM}$ model, but we will assume that $T(t)$ is meaningful at energy scales above that of the CMB, namely $z>1100$, at least up to some limit $T_{\rm max}\approx T_{0}(1+z_{\rm max})$ where $z_{\rm max}\sim e^{1/\widehat{b}}$.
+(See below for explanations of $z_{\rm max}$.)
+Assumption 4.
+Matter and radiation are treated as particle excitations of the Vacuum above its ground state, i.e.
+the Vacuum, with energy density $\rho_{\rm vac}$}.
+We assume $\rho_{\rm vac}$} is the vacuum energy density for all quantum fields, and all of the multiple interactions and parameters of the Standard Model of particle physics are implicit in $\rho_{\rm vac}$}.
+This allows for various deviations of (4) at specific times where various phase transitions can occur, such as the electro-weak phase transition and QCD phase transitions.
+We incorporate matter and radiation after first understanding a universe composed of only vacuum energy, which we will refer to as the dark-universe, and then subsequently add matter and radiation consistent with local energy-momentum conservation $\nabla^{\mu}T_{\mu\nu}=0$, which leads to the usual Friedmann equations with all forms of energy included.
+As already stated, the dark-universe  should be viewed as a skeleton for a complete universe which includes quantum matter and radiation.
+Assumption 5.    We don’t assume gravity is quantized, in part simply because we don’t need it yet to resolve the curvature singularity of the Big Bang, as we will show.
+Under these assumptions all of the QM involved is contained in the QFT computation of $\rho_{\rm vac}$} and it’s renormalization group properties in Minkowski space.
+We are thus implicitly working in the framework of semi-classical gravity, and simply work out the implications of these minimal, well-founded assumptions, to see where it leads us.
+The parameters introduced are $m_{\rm z}$, which sets the scale of the cosmological constant, $\widehat{b}$, and the temperature $T$.
+The Hubble constant $H_{0}$ is just an initial condition for the Friedmann differential equations at the present time $t_{0}$, as is $T_{0}$.
+Taking $\rho_{\rm vac}$} to be its currently measured value in cosmology, namely ${\rho_{\rm vac}}=\Omega_{\Lambda}\rho_{\rm crit}$, where $\Omega_{\Lambda}\approx 0.68$ one finds Planck
+${\rho_{\rm vac}}\approx 5.2\times 10^{-10}\,\dfrac{{\rm Joule}}{{\rm m}^{3}},~~~~~\Longrightarrow~~m_{\rm z}\approx 0.0024\,\mathfrak{g}^{1/4}\,{\rm eV}.$ (5)
+For $\mathfrak{g}={\cal O}(1)$, it is interesting to note that $m_{\rm z}$ is on the order of proposed neutrino masses NeutrinoMasses , however as previously stated, the precise identification of the particle with mass $m_{\rm z}$ will not be necessary for our purposes, and is thus left as an open question.
+Below we will propose additional constraints on the coupling $\mathfrak{g}$.
+From these parameters alone one can develop a pseudo-realistic cosmos which can serve as a skeleton for our own Universe.
+All of the “microscopic” details of the evolution of the Universe obviously depend on the parameters, interactions, and all other bells and whistles of the complete QFT in Minkowski space of the Standard Model of particle physics.
+As we will see, within such a minimal framework one can at least address some of the most profound open problems of modern cosmology: the origin of the Cosmological Constant, the growing Hubble tension, the resolution of $a(t)=0$ singularities associated with the so-called Big Bang, what actually occurred before the Big Bang, why the Universe is flat, and whether current models of inflation play a necessary role or what replaces it.
+This framework may not provide final answers to all of these deep questions, but at the very least it is a novel approach to attempting to address them, and not without some new testable predictions that we will present below.
+Newton’s constant $G_{N}$ is not on the list of basic parameters $m_{\rm z},\widehat{b}$, and $T$ we just presented in the above Assumptions.
+In our proposed framework $G_{N}$ it is not a fundamental constant of Nature, and consequently nor is the Planck length $\ell_{p}=\sqrt{\hbar G_{N}/c^{3}}$.
+From this perspective, natural units should be based on $c,\hbar,{\rm and}~m_{\rm z}$ rather than $c,\hbar,{\rm and}~G_{N}$.
+[^6] From these, one can define a fundamental length $\ell_{\ell}=2\pi\hbar/m_{\rm z}c$.
+Based on the formula (5), if $\mathfrak{g}=1$, then $\ell_{\ell}=5.2$ micrometers, which is huge compared to the Planck length $\ell_{p}$.
+What is more fundamental are mass/energy units.
+Again for $\mathfrak{g}=1$, $m_{\rm z}\approx 0.0024\,{\rm eV}/c^{2}\approx 3.6\times 10^{-36}$   grams, which should be compared with the Planck mass $m_{p}=\sqrt{\hbar c/G_{N}}\approx 2\times 10^{-5}$ grams.
+Based on our current understanding of fundamental particle physics, such as neutrino masses, a fundamental mass of about $0.002\,{\rm eV}/c^{2}$ appears to us more natural than $m_{p}$ which is about the mass of a large amoeba or very fine grain of sand.
+If $G_{N}$ is not a fundamental constant, then what determines its measured value?
 The first observations of Gravity were of course at the surface of the Earth which is manifested as a constant acceleration $g$ for all matter regardless of mass or specific material.
@@
 We keep the fundamental constants $\hbar$ and $c$ explicit in order to make certain conceptual points and to provide some numbers with units.
-“Natural” units based on the Planck length $\ell_{p}$ can be deceptive and are not so natural for this article.
+“Natural” units based on the Planck length $\ell_{p}$ are not so natural for this article.
 In contrast, below we will argue that $G_{N}$ is also environmental, namely
@@
 […]
-Based on the above remarks, what will be central to our exercise is the quantity $\rho_{\rm vac}$}.
-For purposes of illustration, consider a Universe which only consists of a single harmonic oscillator with ground state energy $E_{0}=\hbar\omega/2$, in addition to an imagined observer.
-Since classical mechanics in unaffected by shifts of the potential energy by an arbitrary constant, $E_{0}$ is not measurable in classical physics.
-If all that exists is the harmonic oscillator in empty space, then $E_{0}$ also cannot be measured in a single experiment based on QM.
-There are two known ways to measure $E_{0}$.
-First, one can couple the harmonic oscillator to a generic heat bath at temperature $T$.
-Then one can in principle infer $E_{0}$ from the high temperature behavior of the partition function or its logarithm which is the free energy.
-It’s important that this is based on the free energy and not its derivatives, and thus it is not invariant under shifts of the potential energy.
-The second manner to measure $E_{0}$ is to couple the oscillator to Gravity, since gravity originates from all the matter-energy in the Universe and is sensitive to its ground state energy.
-The fact that the ground state energy can be measured either by coupling to a heat bath or gravity suggests parallels between thermodynamics and gravity that have already been recognized.
-Deceptively simple as it may appear, to study QFT at finite temperature with periodic boundary conditions, one necessarily needs to study quantum fields in curved Euclidean space.
-In 1 spatial dimension the Euclidean space is a torus, and the problem of conformal theory on a torus is a completely solved problem.
-Like Gravity, the Vacuum $|0\rangle$ is omnipresent.
-The Vacuum “knows” the physical properties, such as mass, quantum numbers, and interactions of any particle excitations above it, and is perhaps the only entity that exists everywhere in the Universe and for all times.
-Everything else is some quantum excited state, including people, and typically unstable.
-Nature does not abhor a Vacuum, rather it depends on it.
-[^5] To paraphrase with translation a recognizable philosopher, If you look deeply enough into the Void, it eventually looks back at you.
-So it is reasonable to suppose that the Vacuum, in particular its ground state energy density $\rho_{\rm vac}$}, is at the origin of gravity.
-In other words, if the only way to measure $E_{0}$ is with gravity, then reversing the argument, perhaps quantum vacuum energy is the origin of Gravity.
-For the same reason that one does not necessarily have to quantize the heat bath to study finite temperature QFT, perhaps one does not have to quantize Gravity, however this deep issue will be left as an open question.
-This does not preclude the existence of classical gravitational waves that have been measured, as they can be viewed as classical perturbations of the Vacuum as an elastic medium.
-This article explores this possibility as a theoretical academic exercise to see where it leads.
-This article assumes that $\rho_{\rm vac}$} is well-defined in Minkowski space, namely it is as well defined as the ground state energy of a harmonic oscillator.
-Early considerations of $\rho_{\rm vac}$} based on viewing a free quantized field as a collection of harmonic oscillators with energies $\hbar\omega_{\bf{k}}$, with $\omega_{\bf{k}}=\sqrt{{\bf{k}}^{2}+m^{2}}$ led to the introduction of a cut-off in $|{\bf{k}}|$.
-If this cut-off is taken to be based on the Planck length $\ell_{p}\equiv\sqrt{\hbar G_{N}/c^{3}}$ then $\rho_{\rm vac}$} is much too high by even 100 orders of magnitude and this is perhaps the first formulation of the “cosmological constant problem” Weinberg .
-This naive computation of $\rho_{\rm vac}$} depends on the cut-off and goes to infinity as the cut-off goes to infinity.
-This computation is essentially meaningless, as is the often quoted “worse prediction in physics by $10^{120}$ orders of magnitude”, simply because it is not a real prediction.
-In quantum field theory one is accustomed to dealing with ultra-violet (UV) singularities.
-In recent work ALrhovac1 ; ALrhovac2 we formulated a proper computation of $\rho_{\rm vac}$} which renders it finite and well-defined.
-Since this will be central to subsequent arguments we list its main properties.
-(i)   For 2 dimensional integrable theories $\rho_{\rm vac}$} is exactly computable from the Thermodynamic Bethe Ansatz, and also from the S-matrix bootstrap, and takes the form ${\rho_{\rm vac}}=m_{1}^{2}/2\mathfrak{g}$ where $m_{1}$ is the lightest mass particle.
-In the works ALrhovac1 ; ALrhovac2 this formula was extended to 4 spacetime dimensions:
-${\rho_{\rm vac}}=\dfrac{3}{4}\dfrac{c^{5}}{\hbar^{3}}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}}$ (3)
-where $m_{\rm z}$ is the physical mass of the lightest particle, which we will refer to as the “zeron”, where zero signifies the Vacuum, and $\mathfrak{g}$ is a dimensionless interaction coupling constant.
-Although one does not have the powerful simplifications of integrability in 4 dimensions, this formula was well motivated in ALrhovac1 ; ALrhovac2 and should not be viewed as an ad hoc assumption, although unlike in 2 dimensions it hasn’t yet been fully rigorously proven since the $1/\mathfrak{g}$ dependence indicates it is highly non-perturbative.
-Clearly the above expression is purely quantum mechanical as can be seen from its $\hbar$ dependence.
-The $1/\mathfrak{g}$ dependence can be anticipated, and is in fact necessary, since $\rho_{\rm vac}$} should diverge as $\mathfrak{g}\to 0$ due to an unavoidable logarithmic divergence for free QFT in the UV when the spatial dimension $d$ is evenALrhovac1 .
-[^6] This is essentially another, but refined, version of the “Cosmological Constant Problem”.
-The precise identification of the zeron as a real particle will not be necessary for our arguments, thus this identification will be left as an open question.
-However it was argued in ALrhovac2 that $m_{\rm z}$ is fixed in the deep UV above the electro-weak spontaneous symmetry breaking scale, and that Majorana neutrinos are viable candidates (see (8)).
-[^7] For experimental constraints on neutrino masses see for instance NeutrinoMasses .
-[…]
-${\cal F}=-\dfrac{\pi^{2}T^{4}}{90}\,c(r),~~~~~~{\rm with}~~c(r)=c_{\rm uv}+c_{4}\,r^{4}+\ldots$ (4)
-for some constant coefficients $c_{\rm uv},c_{4}$.
-[^8] In 2 spacetime dimensions $c_{\rm uv}$ is the effective Virasoro central charge.
-Then ${\rho_{\rm vac}}=-c_{4}\pi^{2}m_{\rm z}^{4}/90$.
-This was used in ALrhovac1 to estimate $\rho_{\rm vac}$} for QCD based on finite temperature lattice results.
-(iii)   $\rho_{\rm vac}$} can be independently computed from the form-factor bootstrap that relates the two particle form-factor to the zero particle one, which is a 1-point vacuum expectation value of the field in question ALrhovac2 .
-This is the reason for $m_{\rm z}$ to be the lightest mass particle, since in principle the S-matrix and form factors for other particles can be obtained by bootstrapping the lightest mass particle.
-Applying this to $\langle 0|T_{\mu\nu}|0\rangle$, this leads to a determination of the coupling $\mathfrak{g}$.
-As usual, defining $\cal T$} from the S-matrix $S=1+i{\cal T}$, then
-$\lim_{s\to\infty}{\cal T}(s)=\dfrac{\mathfrak{g}}{m_{\rm z}^{2}}\dfrac{1}{s},$ (5)
-where $s=(p_{1}+p_{2})^{2}$ is the Mandelstam variable for 2-particle scattering.
-In two spacetime dimensions, the two above definitions of $\rho_{\rm vac}$} were shown to agree exactly for a variety of integrable quantum field theories ALrhovac1 ; ALrhovac2 .
-As previously mentioned, in one spatial dimension it takes the form ${\rho_{\rm vac}}=c^{3}m_{1}^{2}/2\hbar\mathfrak{g}$ where $m_{1}$ is the lightest particle and $\mathfrak{g}$ can be explicitly computed.
-Various consistency checks on this formula were made, including that it must vanish for supersymmetric theories and also theories with a fractional supersymmetries where the hamiltonian is in the center of the universal enveloping algebra of the conserved charges.
-(See ALrhovac1 and references therein.) The coupling $\mathfrak{g}$ can be positive or negative, and $\rho_{\rm vac}$} can even oscillate in sign for a given model, for instance for the sine-Gordon model where it oscillates from positive to negative, where it is zero at the fractional supersymmetric points.
-In 4 spacetime dimensions it was not rigorously shown that equations (4), (5) lead to the same formula for $\rho_{\rm vac}$}, thus (3) should still be viewed as well-motivated but still conjectural.
-For the remainder of the article we simply assume the well-motivated formula (3).
-Some independent support for the result (3) comes from Swampland ideas Festina1 ; Festina2 . By studying a charged particle of mass $m$ in the presence of a black hole, it was argued that
-${\rho_{\rm vac}}<\dfrac{m^{4}}{2e^{2}}$ (6)
-where here $\mathfrak{g}$ is identified with the quantum electrodynamic (QED) fine structure constant $\alpha=e^{2}/\hbar c$.
-Actual QED will not play a role in this article since we assume the zeron is electrically neutral, however, as we will see, it will be instructive to compare couplings and renormalization group parameters with QED, just to check if they are reasonable.
-Our result (3) is stronger than (6) since it is not an inequality, but it is still consistent with (6).
-As previously stated, in this article we do not propose what the hypothetical zeron particle actually is, nor even that it is a detectable particle, however we can constrain it’s mass $m_{\rm z}$ from the cosmologically measured value of $\rho_{\rm vac}$}.
-Taking it to be ${\rho_{\rm vac}}=\Omega_{\Lambda}\rho_{\rm crit}$, where $\Omega_{\Lambda}\approx 0.68$ one finds Planck
-${\rho_{\rm vac}}=5.2\times 10^{-10}\,\dfrac{{\rm Joule}}{{\rm m}^{3}}.$ (7)
-Using this value, then (3) implies
-$m_{\rm z}=0.0024\,\mathfrak{g}^{1/4}\,{\rm eV}.$ (8)
-For $\mathfrak{g}={\cal O}(1)$, it is interesting to note that $m_{\rm z}$ is on the order of proposed neutrino masses NeutrinoMasses .
-Below we will propose additional constraints on the coupling $\mathfrak{g}$.
 It is instructive to compare the perspective on Gravity developed in this article with some proposed theories of quantum gravity, since this clearly delineates some important distinctions.
-We take “quantum gravity” to mean a theory where the space-time metric $g_{\mu\nu}(x)$ itself is quantized which implies the existence of spin 2 gravitons.
+We take “quantum gravity” to mean a theory where the space-time metric $g_{\mu\nu}(x)$ itself is quantized which implies the existence of massless spin 2 gravitons.
+Based on this delineation, quantum gravity plays no role in the present article, and in spite of this our model avoids the usual curvature singularity when the scale fact $a(t)=0$.
 Although the definition of the Planck length $\ell_{p}$, defined from $G_{N}$ as $G_{N}=c^{3}\,\ell_{p}^{2}/\hbar$, is a priori just based on dimensional analysis, one expects that $\ell_{p}$ is a relevant scale when considering graviton-graviton scattering.
-The latter does not play a role in this article, which makes one question whether $\ell_{p}$ is a fundamental constant, however this issue will not be completely settled in the present cosmological context.
-[^9] Based on this article, units based on $\hbar$, $c$ and $m_{\rm z}$ are more natural.
-A significant result of string theory is that it provides a consistent prescription for the computation of graviton scattering that is finite StringTheory .
+A significant result of string theory, which continues to be its main motivation, is that it provides a consistent prescription for the computation of graviton scattering that is finite StringTheory .
 Thus gravity “emerges” from string theory, in the sense that Newton’s constant $G_{N}$ is determined by matching graviton scattering with low energy gravity.
-This leads to a formula for $G_{N}$:
+This leads to a formula for $G_{N}$ which is completely determined by physics at UV energy scales:
 $G_{N}\sim c_{Y}\,g_{s}^{2}\,\,\dfrac{c^{3}\ell_{s}^{2}}{\hbar}$ (7)
 […]
-With the above introductory remarks, let us summarize the remainder of this article.
-The two primary assumptions we make are the formula (3) for $\rho_{\rm vac}$} which was well-motivated by the works ALrhovac1 ; ALrhovac2 , and the renormalization group flow for $\mathfrak{g}$ in equation (35) below.
-In the next section we formulate a gravitational version of the electromagnetic Casimir effect, and provide an argument for how gravity can arise from $\rho_{\rm vac}$}.
-In particular we show how Einstein’s field equations emerge, in this context in the form of Friedmann’s equations.
-The formula (2) leads to new interpretation of the Gibbons-Hawking entropy of de Sitter space, wherein the quantum informational “bits” are identified as quantized massless particles at the horizon with wavelength $\lambda=2\pi{{\rm R}_{\scriptscriptstyle\infty}}$.
-The analysis in Section II is independent of the form $\rho_{\rm vac}$} in (3), and we turn to understanding the implications of (3) in Section III.
-There we assume $\mathfrak{g}$ is a marginal coupling with a renormalization group (RG) flow, which as we will argue is well-motivated, and show that this leads to an effective, induced RG flow for Newton’s constant $G_{N}$ as a function of an energy scale $\mu$.
+With the above introductory remarks, let us summarize the remainder of this article as outlined in the Table of Contents.
+In the next section we formulate a gravitational version of the electromagnetic Casimir effect, and provide a heuristic argument for how gravity can arise from $\rho_{\rm vac}$}.
+In particular we show how Einstein’s field equations emerge, in this context in the form of Friedmann’s equations where $G_{N}$ is given by (6).
+The formula (6) leads to new interpretation of the Gibbons-Hawking entropy of de Sitter space, wherein the quantum informational “bits” are identified as quantized massless particles like photons at the horizon with wavelength $\lambda=2\pi{{\rm R}_{\scriptscriptstyle\infty}}$, as described in Section III.
+The analysis in Section II is independent of the formula for $\rho_{\rm vac}$} in (2), and we turn to understanding the implications of it in Section IV.
+There we argue that the RG flow for the coupling $\mathfrak{g}$ induces an effective RG flow for Newton’s constant $G_{N}$ as a function of an energy scale $\mu$.
 [^10] We previously published some preliminary work on the possible implications of the RG flow of $\mathfrak{g}$ for cosmology in LeClairUniverse , however the present article supersedes it.
 As is standard in Cosmology, we tie this energy scale $\mu$ to the redshift $z$ and obtain equation (55).
-Solving the resulting equations for the scale factor $a(t)$ results in a significant modification of de Sitter space that is essentially an inverted gaussian and at no time is $a(t)=0$, which signifies no classical type of geometric singularity since for all times $a(t)>a_{\rm min}$.
+In Section V we solve exactly the modified Friedmann equations for the dark-universe, which results in a significant modification of de Sitter space that is essentially an inverted gaussian and at no time is $a(t)=0$, which signifies no classical type of geometric curvature singularity since for all times $a(t)>a_{\rm min}$.
 The solution to the scale factor satisfies an interesting symmetry between the far past and far future, namely $a(t)=a(-t+2t_{\rm min})$ where $a(t_{\rm min})=a_{\rm min}$.
-There is no Big Bang in our model since the solutions in past and future match at the time $t_{\rm min}$.
-Extending our solution to the far past, the time evolution of the Universe is more analogous to the swing of a pendulum.
-Our interpretation of this is that the age of the Universe is actually infinite since the solution to $a(t)$ extends to the far past without ever reaching the singularity at $a=0$.
-In Section IV we add matter and radiation, and due to the RG flow of $G_{N}$ in (43) there is again no $a=0$ singularity and the symmetry between the past and future continues to hold.
-This implies very small logarithmic corrections to the $\Lambda{\rm CDM}$ model if $\widehat{b}$ is small.
+There is no Big Bang normally associated with the singularity $a(t)=0$, rather at the time $t_{\rm min}$ the universe is at its hottest, and is more properly referred to as a hot Big Bang.
+In our model since the solutions in past and future match at the time $t_{\rm min}$, one can address what happened before the Big Bang.
+Extending our solution to the far past, the time evolution of our model universe is more analogous to the harmonic swing of a pendulum.
+Our interpretation of this is that the age of this model universe is actually infinite since the solution to $a(t)$ extends to the far past without ever reaching the singularity at $a=0$.
+In Section VI we add matter and radiation, and due to the RG flow of $G_{N}$ in (55) there is again no $a=0$ singularity and the symmetry between the past and future continues to hold.
+This implies very small logarithmic corrections to the $\Lambda{\rm CDM}$ model if $\widehat{b}\equiv\,b\,\mathfrak{g}_{0}/2\pi$ is small.
 We apply these ideas to the so-called Hubble tension, wherein the discrepancy of measurements of Hubble constant $H_{0}$ is proposed to arise from the fact that the two contradictory measurements rely on data that refers to different epochs where Newton’s constant is effectively different since the energy scales involved are different.
+In Section VII we perform some checks of our model by comparing with a variety of cosmological and other types of experimental data.
+The strongest support comes from recently observed trends in the Hubble tension for low-z supernovae.
 II Emergence of Einstein’s equations from a Gravitational Casimir effect
 In this section we formulate a gravitational version of the electromagnetic Casimir effect.
+These considerations are independent of the specific QFT details that lead to the formula (2).
+In this section we only assume that it is well-defined and finite, and will turn to the implications of the formula (2) in the next section.
 Without assuming Einstein’s field equations, we present heuristic arguments for how they emerge in this context.
-These arguments lead to the formula (2) for Newton’s constant $G_{N}$ expressed in terms of length and energy scales in the Universe.
-This doesn’t rely on the explicit formula (3) for $\rho_{\rm vac}$}, where its implications will be studied in the next section.
+These arguments lead to the formula (6) for Newton’s constant $G_{N}$ expressed in terms of length and energy scales in our model universe.
+The discussion of this section is not absolutely necessary for understanding the remainder of this article.
+The reader may prefer to simply assume Einstein’s field equations and move onto the next section where QFT properties of $\rho_{\rm vac}$} are invoked.
 Let us first review the electromagnetic Casimir effect, which is well understood theoretically and has been measured Casimir .
 This Casimir effect requires two uncharged, parallel conducting plates in vacuum separated by a distance $L$.
+The plates are required to radiate and confine photons between the plates.
 The energy density of quantized photons between the plates $\rho_{\rm rad}$ leads to a measurable force between the plates $F=\partial_{L}E$ where $E=\rho_{\rm rad}\cdot V$ where $V$ is the volume between the plates.
 It is important to point out that the theoretical predictions, which are confirmed in the measurements, do not measure the vacuum energy density $\rho_{\rm vac}$} but rather the energy density of free photons to leading order, whose pressure is $p=w\rho_{\rm rad}$, where $w=1/3$, whereas genuine vacuum energy density has w=1w=-1.
-In fact, since the free theory of photons is conformally invariant, it’s vacuum energy density ${\rho_{\rm vac}}=0$.
-II.1 Formulating a gravitational Casimir effect without plates
-For gravity, there is no well-formulated analog of the electromagnetic Casimir effect, since hypothetical gravitons cannot be constrained by material boundaries like plates.
-In any case, if any experiment could be imagined that overcomes these obvious difficulties, its effects would be much too small to be measurable.
-This led us to formulate a gravitational Casimir effect without material boundaries, namely without plates, that is sensitive to $\rho_{\rm vac}$}, where the latter is the vacuum energy density of all quantum matter and radiation excluding gravity.
+In fact, since the theory of non-interacting photons is conformally invariant, it’s vacuum energy density ${\rho_{\rm vac}}=0$ in the leading conformal limit.
+For Gravity, there is no well-formulated and strict analog of the electromagnetic Casimir effect, since gravitational waves and hypothetical gravitons cannot be constrained by material boundaries like conducting plates.
+In any case, if any experiment could be imagined that overcomes these obvious difficulties, its effects would be much too small to be measurable if they originate from gravitons.
+This leads us to formulate a gravitational Casimir effect without material boundaries, namely without plates, that is sensitive to $\rho_{\rm vac}$}, where the latter is the vacuum energy density of all quantum matter and radiation excluding gravity.
 Throughout this paper we assume ${\rho_{\rm vac}}>0$ such that we are dealing with de Sitter space.
-We make three additional assumptions: (i) $\rho_{\rm vac}$} is well-defined and finite in flat Minkowski space for the entirety of quantum fields for matter and radiation without gravity, defined in (1).
-This quantity $\rho_{\rm vac}$} is in principle computable without gravity, as explained in the Introduction.
-Secondly, (ii) The Vacuum has no time dependence originating from Einstein’s equations, namely $\partial_{t}{\rho_{\rm vac}}=0$.
+Our working assumptions were presented in the Introduction.
+One additional assumption is that the vacuum has no time dependence originating directly from the time dependence of the resulting Einstein’s equations, namely $\partial_{t}{\rho_{\rm vac}}=0$.
 However as we will see in the next section, tying the RG energy scale to the redshift $z$ leads to an additional induced time dependence.
-Third, (iii) There is no additional and independent free parameter $\Lambda$ in Einstein’s equations corresponding to a classical cosmological constant.
-This is consistent with the fact that $\rho_{\rm vac}$} is not invariant under arbitrary shifts of the potential energy since it is in principle computable from the logarithm of the partition function and not its derivatives.
-We carry out this construction in arbitrary spatial dimension $d$ in order to make certain points.
-Consider first a relativistic QFT in $d=1$ spatial dimension, from which one can compute $\rho_{\rm vac}$}.
-There are an infinite number of integrable QFT’s where one can calculate $\rho_{\rm vac}$} exactly; see for instance the examples in ALrhovac1 ; ALrhovac2 .
+We carry out this construction in arbitrary spatial dimension $d$ in order to make certain conceptual points.
+Let us begin with only $1$ more spatial dimension than the harmonic oscillator in QM discussed in the Introduction which corresponds to $d=0$.
+Thus, consider a relativistic QFT in $d=1$ spatial dimension, from which one can compute $\rho_{\rm vac}$}.
+There are an infinite number of integrable QFT’s for $d=1$ where one can calculate $\rho_{\rm vac}$} exactly; see for instance the examples in ALrhovac1 ; ALrhovac2 which are reviewed in Section IVA.
 Let the spatial dimension be a finite size segment of length $2R$.
 Since total vacuum energy is $E=2R\,{\rho_{\rm vac}}$, there is a force $F=dE/dR=2{\rho_{\rm vac}}$ that can in principle be measured by moving the “walls”, in this case simply the endpoints of the segment.
-However if the endpoints do not consist of matter being held in place, the effect of this force is expansion or contraction of the size of the segment, depending on the sign of $\rho_{\rm vac}$}.
-This naturally leads us to introduce geometry, namely a metric $g_{\mu\nu}$:
+However if the endpoints do not consist of matter being held in place, that is to say there is no analog of conducting plates, the effect of this force is expansion or contraction of the size of the segment, depending on the sign of $\rho_{\rm vac}$}.
+This naturally leads us to introduce geometry, namely a metric $g_{\mu\nu}$ which can incorporate this spatial expansion:
 $ds^{2}=-g_{\mu\nu}(x)dx^{\mu}dx^{\nu}=-c^{2}dt^{2}+a(t)^{2}dx^{2}.$ (8)
-The dimensionless scale factor can be thought of as $a(t)=R(t)/R_{0}$ where $R_{0}$ is the scale factor today, such that at the present time $t_{0}$, $a(t_{0})=1$.
+The hypothetical experimental setup is the Universe itself, since there are no conducting plates involved, and as we will argue, the cosmic horizon plays the role of such plates.
+The dimensionless scale factor by definition satisfies $a(t_{0})=1$ at the present time $t_{0}$.
 In order to make sense of Newton’s 2nd law, one needs to introduce an inertial mass $M$.
@@
 $MR\,\ddot{a}=2a\,{\rho_{\rm vac}}~~~\Longrightarrow~~\dfrac{\ddot{a}}{a}=\dfrac{2{\rho_{\rm vac}}}{MR}.$ (9)
-We next require $\partial_{t}{\rho_{\rm vac}}=0$. One needs another equation to enforce this, which is $\ddot{a}/a=(\dot{a}/a)^{2}$:
+We next require $\partial_{t}{\rho_{\rm vac}}=0$. One needs another equation to enforce this, which is just $\ddot{a}/a=(\dot{a}/a)^{2}$:
 $\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{2{\rho_{\rm vac}}}{MR}.$ (10)
 One can easily check that one time derivative of (10) combined with (9) implies $\partial_{t}{\rho_{\rm vac}}=0$.
-Let us extend the above arguments to $d$ spatial dimensions, where now $R$ is a radius. Define $V_{d}(R)$ to be the volume of a sphere in $d$ dimensions. Then one has
-$MR\,\ddot{a}=\dfrac{a}{d}\,\dfrac{\partial V_{d}(R)}{\partial R}{\rho_{\rm vac}},~~~~~V_{d}(R)=\dfrac{2\pi^{d/2}R^{d}}{d\,\Gamma(d/2)}.$ (13)
-The factor of $1/d$ on the right hand side of the first equation is due to the vectorial nature of the force. To summarize, we have two equations after imposing $\partial_{t}{\rho_{\rm vac}}=0$:
+Let us extend the above arguments to $d$ spatial dimensions, and assume rotational symmetry, where now $R$ is a radius. Define $V_{d}(R)$ to be the volume of a sphere in $d$ dimensions. Then one has
+$MR\,\ddot{a}=\dfrac{a}{d}\,\dfrac{\partial V_{d}(R)}{\partial R}{\rho_{\rm vac}},~~~~~V_{d}(R)=\dfrac{2\pi^{d/2}R^{d}}{d\,\Gamma(d/2)},$ (11)
+where $\Gamma$ is the usual $\Gamma$-function.
+The factor of $1/d$ on the right hand side of the first equation is due to the vectorial nature of the force.
+To summarize, we have two equations after imposing $\partial_{t}{\rho_{\rm vac}}=0$:
 $\dfrac{\ddot{a}}{a}=\dfrac{R^{d-2}}{M}\left(\dfrac{2\pi^{d/2}}{d\,\Gamma(d/2)}\right){\rho_{\rm vac}}$ (12)
@@
 $G_{N}$ can be identified by comparing with Einstein’s field equations for this particular metric.
-We adopt standard conventions for the Einstein-Hilbert action in $d$ dimensions:
+We adopt standard conventions for the Einstein-Hilbert action in $d$ spatial dimensions:
 ${\cal S}=\dfrac{1}{16\pi G_{d}}\int d^{d+1}x\,\sqrt{-g}\,\left({\cal R}+{\cal L}_{\rm matter}\right).$ (15)
 The metric is as in (8), with with $dx^{2}$ replaced with $d\vec{x}\cdot d\vec{x}$.
-There is no a priori reason to include a curvature term proportional the standard $k\in\{-1,0,1\}$ in the spatial metric based on the above formulation of the Casimir effect.
+Based on the above formulation of the gravitational Casimir effect, there is no a priori reason to include a curvature term in the spatial part of the metric proportional the standard $k\in\{-1,0,1\}$.
 In other words $k=0$ is natural for our formulation, since, from our perspective, Gravity originates from $\rho_{\rm vac}$} in Minkowski space, and the latter clearly has $k=0$.
@@
 ${\cal R}_{\mu\nu}-\tfrac{1}{2}g_{\mu\nu}\,{\cal R}=\dfrac{8\pi G_{d}}{c^{d+1}}\,T_{\mu\nu}.$ (16)
+It’s important to note that there is no room, i.e.
+no a priori reason, to introduce an additional classical free parameter $\Lambda$ corresponding to an independent cosmological constant term in Einstein’s equations based on the above arguments.
+Furthermore, we already explained that once the QFT is decided upon, there is no room for arbitrary shifts in $\rho_{\rm vac}$} since it is based on the logarithm of its free energy density and not its derivatives, and this avoids the fine-tuning issues normally associated with a variable free parameter $\Lambda$ corresponding to a cosmological constant.
 If one carefully keeps track of the $d$-dependence, this leads to the following Friedmann equations [^11] The $d$ dependence of this formula should be a well-known result.
-One can find it for instance in Gibbons .
+One can find it for instance in GibbonsDdim .
 $\dfrac{\ddot{a}}{a}=-\dfrac{8\pi G_{d}}{d(d-1)c^{2}}\left[(d-2)\rho+d\,p\right]$ (17)
@@
 $R=R_{H}=\dfrac{c}{H},$ (20)
-which is referred as the Hubble radius. With the identification of $G_{d}$ in (21), the equation (20) is equivalent to
+which is referred as the Hubble radius. With the identification of $G_{d}$ in (19), the equation (18) is equivalent to equation (14):
 ${\rho_{\rm vac}}V_{d}(R_{H})=Mc^{2}.$ (21)
@@
 First of all, we have not yet included matter and radiation as excitations over the Vacuum, where $H$ is no longer a constant, and we will return to this below.
-Let us simply point out that if one takes the experimental value (7) and identifies $H=\sqrt{\Omega_{\Lambda}}H_{0}$ where $H_{0}$ is Hubble constant as measured by the CMB (52) and $\Omega_{\Lambda}=0.68$, then the above expression evaluates to $G_{N}=6.72\times 10^{-11}$ ${\rm m}^{3}/{\rm kg}\,{\rm s}^{2}$.
+Let us just point out that if one takes the experimental value (5) and identifies $H=\sqrt{\Omega_{\Lambda}}H_{0}$ where $H_{0}$ is Hubble constant as measured by the CMB (64) and $\Omega_{\Lambda}=0.68$, then the above expression evaluates to $G_{N}=6.72\times 10^{-11}$ ${\rm m}^{3}/{\rm kg}\,{\rm s}^{2}$.
 Let us note that the same computation using $H_{0;{\rm SN}}$ in (64) yields a higher value for $G_{N}$.
 This will be addressed below in connection with the “Hubble tension” by taking into account the explicit expression (2) for $\rho_{\rm vac}$}.
-II.2 An interpretation of the Gibbons-Hawking entropy of de Sitter space
+III A reinterpretation of the entropy of de Sitter space and black holes
 There are some important results in semi-classical gravity where the gravitational field is not quantized, namely they do not rely on the existence of gravitons, but the result still depends on the Planck length $\ell_{p}$, such as the Bekenstein-Hawking entropy and temperature for black holes Bekenstein ; Hawking .
 In this present context, the analog is the Gibbons-Hawking entropy for de Sitter space GibbonsHawking .
-Since these results depend on $G_{N}$, the expression (2) for $G_{N}$ leads to an interesting new interpretation of this entropy.
+Since we have brought into scrutiny whether $G_{N}$ and thus Planck length $\ell_{p}$ are fundamental parameters, this led us to re-interpret these entropies, since they are expressed in terms of $\ell_{p}$.
+The expression (6) for $G_{N}$ leads to such a new interpretation that is very simple and illuminating.
 […]
 $S=k_{B}N_{\rm bits}.$ (25)
-The Gibbons-Hawking entropy was thus proposed to be
+The Gibbons-Hawking entropy $S_{GH}$ was proposed to be
 $S_{GH}\equiv\dfrac{k_{B}}{4}\dfrac{A}{\ell_{p}^{2}}=k_{B}\left(\dfrac{\pi c^{3}}{\hbar}\right)\dfrac{{{\rm R}^{2}_{\scriptscriptstyle\infty}}}{G_{N}}.$ (26)
@@
 Here each bit is commonly interpreted as a pixelation of the area $A$ into quantized units of area $4\ell_{p}^{2}$, suggesting that spacetime itself is somehow quantized into these minimal area units at the horizon.
+The overall factor of $1/4$ is the hardest to explain in the above formula.
 Such a quantization of spacetime itself remains a somewhat vague notion, and difficult to make sense of mathematically.
@@
 […]
-Let us point out that the above interpretation of the Gibbons-Hawking entropy also applies to the Bekenstein-Hawking entropy of a black hole.
+The above reinterpretation of the Gibbons-Hawking entropy also applies to the Bekenstein-Hawking entropy of a black hole, where again the bits are quantized massless particles.
+[^14] For super massive black holes, the quantization of spacetime itself into bits at the super large horizon seems to us as rather implausible.
 The reason is that (6) is the correct relation between the mass $M_{\bullet}$ of the black hole and its radius $R_{\bullet}$ if one makes the replacement ${{\rm M}_{\scriptscriptstyle\infty}},{{\rm R}_{\scriptscriptstyle\infty}}\to M_{\bullet},R_{\bullet}$:
 $G_{N}=\dfrac{c^{2}}{2}\dfrac{R_{\bullet}}{M_{\bullet}}.$ (30)
-From this perspective, $G_{N}$ is a universal constant which is the same for every black hole in the Universe. Repeating the above arguments for the Bekenstein-Hawking entropy,
+From this perspective, $G_{N}$ is a constant which is the same for every black hole in the Universe. Repeating the above arguments for the Bekenstein-Hawking entropy,
 $S_{BH}=\dfrac{k_{B}}{4}\dfrac{A}{\ell_{p}^{2}}=k_{B}\left(\dfrac{\pi c^{3}}{\hbar}\right)\dfrac{R_{\bullet}^{2}}{G_{N}}=k_{B}\,\dfrac{E_{\bullet}}{E_{\lambda}},~~~~~~{\rm with}~E_{\bullet}=M_{\bullet}c^{2}~~{\rm and}~~E_{\lambda}=\dfrac{\hbar c}{2\pi R_{\bullet}},$ (31)
 where we have used (30) in the second equation. The Hawking temperature $T_{H}$ of the black hole is again given by (29), namely $k_{B}T_{H}=E_{\lambda}$.
-III Induced renormalization group flow for Newton’s constant
-III.1 Effective RG flow for Newton’s constant based on the RG flow for $\rho_{\rm vac}$}.
-The results of the last section apply whatever $\rho_{\rm vac}$} is.
-In this section we use the expression (3) to obtain further results.
+IV Induced renormalization group flow for Newton’s constant
+IV.1 Justifying our formula for $\rho_{\rm vac}$}.
+In this subsection we review the analysis that led us to propose the formula (2) for $\rho_{\rm vac}$} in ALrhovac1 ; ALrhovac2 .
+As motivation, let us simply consider a real scalar field $\phi$ in $d+1$ spacetime dimensions with euclidean action
+${\cal S}=\int d^{d+1}x\,\left(\tfrac{1}{2}\partial_{\mu}\phi\partial_{\mu}\phi+V(\phi)\right).$ (32)
+Suppose the potential is simply a mass term $V(\phi)=m\phi^{2}/2$ such that spectrum consists of particles of energy $\omega_{\bf{k}}=\sqrt{{\bf{k}}^{2}+m^{2}}$.
+Treating the free quantized field as a collection of harmonic oscillators with energy $\hbar\omega_{\bf{k}}/2$, naively the vacuum energy density is
+$\rho_{\rm vac,cutoff}=\dfrac{s}{2}\int_{|{\bf{k}}|\leq k_{c}}\dfrac{d^{d}{\bf{k}}}{(2\pi)^{d}}\,\sqrt{{\bf{k}}^{2}+m^{2}},$ (33)
+where $s=+/$- corresponds to bosons/fermions. Since the above integral is divergent, we have introduced a cut-off $k_{c}$. The leading terms for $k_{c}\gg m$ are
+$\rho_{\rm vac,cutoff}=s\begin{cases}\dfrac{k_{c}^{2}}{4\pi}+\dfrac{m^{2}}{4\pi}\,\log(2k_{c}/m)~~~~~~~~~(d=1)\\ \\ \dfrac{k_{c}^{3}}{12\pi}-\dfrac{m^{3}}{12\pi}~~~~~~~~~~~~~~~~~~~~~~~(d=2)\\ \\ \dfrac{k_{c}^{4}}{16\pi^{2}}-\dfrac{m^{4}}{32\pi^{2}}\,\log(2k_{c}/m)~~~~~(d=3)\end{cases}$ (34)
+The above calculation is what led Weinberg to state a Cosmological Constant Problem Weinberg .
+Namely, if $k_{c}$ is taken to be the Planck scale, then the leading term is off by $120$ orders of magnitude.
+There are two obvious problems with this calculation.
+First of all, there is no justification for $k_{c}$ to be the Planck scale since we are dealing with a free quantum field theory in Minkowski space with no gravity.
+Secondly, one is accustomed to ultra-violet divergences in QFT and how to deal with them in order to compute physical quantities.
+In such a renormalization procedure the leading $k_{c}^{d+1}$ is discarded.
+However note that for $d+1$ even there is an unavoidable $\log k_{c}$ divergence, unlike $d+1$ odd, and further renormalization is required in order to obtain something finite and independent of the cut-off.
+However if there are no additional interactions in $V(\phi)$, there is no clear and physical way to remove the $\log k_{c}$ divergence.
+Based on the above equations (34), the issue of how to deal with the remaining $\log k_{c}$ divergence is essentially identical in $d=1$ and $d=3$ spatial dimensions, thus one can gain insights on the problem from interacting QFT’s in $d=1$ that are exactly solvable, i.e.
+integrable.
+Firstly, one can argue that in $d+1$ spacetime dimensions ${\rho_{\rm vac}}\propto m^{d+1}/g$ where $m$ is a physical mass, and $g$ is an interaction coupling, such that $\rho_{\rm vac}$} diverges as $g\to 0$ reflecting the unavoidable divergence in the free theory displayed in (34).
+If $m$ is a physical mass, then by dimensional analysis, $g$ is a dimensionless coupling for an interacting theory.
+That ${\rho_{\rm vac}}\propto 1/g$ implies that it is intrinsically non-perturbative.
+It turns out this is exactly what occurs for integrable QFT in 2 spacetime dimensions ZamoZamo0 ; ZamoTBA .
+Namely for all these integrable theories,
+${\rho_{\rm vac}}=\dfrac{m_{1}^{2}}{2\mathfrak{g}}~~~~~~~~~~~~~~~~~~~~~~~(d=1)$ (35)
+where $m_{1}$ is the mass of the lightest massive particle, and $\mathfrak{g}$ represents a dimensionless coupling, which is a non-perturbative function of the couplings $g$ in $V(\phi)$.
+The formula (35) can be derived from the thermodynamic Bethe ansatz, which was first formulated by Yang and Yang YangYang , and generalized to relativistic theories by Al.
+Zamolodchikov ZamoTBA .
+That $m_{1}$ is the mass of the lightest massive particle follows from the fact that any particle can probe the vacuum to determine $\rho_{\rm vac}$} and the result should be the same, combined with the S-matrix bootstrap which implies that in principle the S-matrix for higher mass particles can be obtained by bootstrapping the lightest mass particle.
+For integrable theories all of this is well-understood and previously worked out explicitly for many examples.
+See for instance KlassenMelzer , and Corrigan for the affine Toda theories based on an arbitrary Lie group.
+Let us illustrate with the so-called sinh-Gordon model, which is the affine Toda theory based on the Lie group SU(2), with potential
+$V(\phi)=2\mu\cosh(\sqrt{8\pi}\,b\,\phi),$ (36)
+where $\mu$ sets the energy scale for the mass $m$ of the single particle in the spectrum, and $b$ is a dimensionless coupling.
+[^15] Not to be confused with $b$ in (3).
+Here $\mathfrak{g}=-4\pi\sin(\pi\gamma),~~\gamma\equiv b^{2}/(1+b^{2})$, and one finds
+${\rho_{\rm vac}}=-\dfrac{m^{2}}{8\pi b^{2}}\left(1+b^{2}+\dfrac{\pi^{2}}{6}b^{4}+\ldots\right).$ (37)
+The above result can also be derived in Feynman diagram perturbation theory DestrideVega .
+The leading term is non-perturbative in the coupling $b$ and the terms in parentheses correspond to perturbative corrections.
+Indeed one recognizes $\zeta(2)=\pi^{2}/6$ which is commonplace in perturbation theory, and one can easily check the higher orders in $b$ involve $\zeta(2n)$ for integer $n$.
+As a warm-up exercise for 4 spacetime dimensions, it is instructive to understand the leading $1/b^{2}$ term in (37) without relying on integrability.
+To this end we expand the $\cosh$ potential in (35) in powers of $\phi$:
+$V(\phi)=2\mu+\dfrac{m^{2}}{2}\phi^{2}+\dfrac{\lambda}{4!}\phi^{4}+{\cal O}(\phi^{6}),~~~~{\rm where}~~~m^{2}=16\pi\mu b^{2},~~~\lambda=128\mu\pi^{2}b^{4}\,.$ (38)
+Minimizing the potential $\partial_{\phi}V(\phi)=0$, in addition to a minimum at $\phi=0$ there is a minimum for non-zero $\phi=\overline{\phi}$ where $\overline{\phi}^{2}=-6m^{2}/\lambda$.
+Ignoring for the moment that this $\overline{\phi}$ is only real for negative $\lambda$, one finds
+${\rho_{\rm vac}}=V(\overline{\phi})=2\mu-\dfrac{3}{2}\dfrac{m^{4}}{\lambda}=-\mu=-\dfrac{m^{2}}{16\pi b^{2}}\,$, (39)
+where we have used the expressions for $m^{2}$ and $\lambda$ in (38).
+This agrees with (37) up to a factor of $2$, the discrepancy arising from the fact that we truncated the $\cosh$ potential to order $\phi^{4}$.
+One can obtain the exact leading term with a similar semi-classical calculation starting from the $\cosh$ potential (36), where the minimum is at $\overline{\phi}=0$ such that $V(\overline{\phi})=2\mu$.
+One delicate point is that in the thermodynamic Bethe ansatz the scalar particle corresponding to the field $\phi$ must be treated as a fermion since in the high energy limit the S-matrix S=1S=-1.
+Thus ${\rho_{\rm vac}}=-2\mu$ which equals the leading term in (37).
+It is important to note that the value of $\rho_{\rm vac}$} in (39) does not require spontaneous symmetry breaking (SSB).
+Let us now turn to 4 spacetime dimensions, where one does not have the powerful tools of integrability available.
+In recent work ALrhovac1 ; ALrhovac2 we formulated a proper computation of $\rho_{\rm vac}$} which renders it finite and well-defined in 4 dimensional Minkowski space and its theoretical basis is the same as in two spacetime dimensions.
+In ALrhovac1 we showed how to calculate $\rho_{\rm vac}$} from QFT at finite temperature $T$, which parallels the its computation from the thermodynamic Bethe ansatz in $d=1$.
+Riemann’s zeta function $\zeta(s)$ and its fundamental properties such as the functional equation that relates $\zeta(s)$ to $\zeta(1-s)$ played an essential role in the analysis in ALrhovac1 , and is analogous to modularity in $d=1$.
+Namely, let $m_{\rm z}$ be the fundamental renormalized scale of physical particle masses where $m_{\rm z}$ represents the lightest mass particle, and $c(r)$ where $r=m_{\rm z}c^{2}/k_{B}T$ be the scaling function such that the free energy density is given by
+${\cal F}=-\dfrac{\pi^{2}T^{4}}{90}\,c(r),~~~~~~{\rm with}~~c(r)=c_{\rm uv}+c_{4}\,r^{4}+\ldots$ (40)
+for some constant coefficients $c_{\rm uv},c_{4}$.
+[^16] In 2 spacetime dimensions $c_{\rm uv}$ is the effective Virasoro central charge.
+In 4 spacetime dimensions, $c_{\rm uv}$ is known from conformal field theory if the theory is asymptotically free, as in QCD.
+Then
+${\rho_{\rm vac}}=-c_{4}\pi^{2}m_{\rm z}^{4}/90.$ (41)
+This was used in ALrhovac1 to estimate $\rho_{\rm vac}$} for QCD based on finite temperature lattice results where $c_{\rm uv}$ is known since QCD is asymptotically a free conformal field theory.
+One can also in principle compute $\rho_{\rm vac}$} directly from the S-matrix at zero temperature using the form-factor bootstrap ALrhovac2 .
+One advantage of this formulation that all masses are the physical (renormalized) ones once the 2-particle form factor is properly normalized.
+The form-factor bootstrap relates the two particle form-factor to the zero particle one, which is a 1-point vacuum expectation value of the field in question.
+As in two spacetime dimensions, we take $m_{\rm z}$ to be the physical mass of the lightest particle, since in principle the S-matrix and form factors for other particles can be obtained by bootstrapping this lightest mass particle.
+This is where we need Assumption 1 stated in the Introduction, namely that all particle excitations are coupled.
+In connection with this, in ALrhovac2 we stated the principle of particular democracy, wherein any particle can be used to probe the Vacuum and its energy density, and they should give the same value for $\rho_{\rm vac}$}, and this is known to be true for the $d=1$ models considered there.
+Applying this to $\langle 0|T_{\mu\nu}|0\rangle$, this leads to a determination of the coupling $\mathfrak{g}$.
+As usual, defining $\cal T$} from the S-matrix $S=1+i{\cal T}$, then
+$\lim_{s\to\infty}{\cal T}(s)=\dfrac{\mathfrak{g}}{m_{\rm z}^{2}}\dfrac{1}{s},$ (42)
+where $s=(p_{1}+p_{2})^{2}$ is the Mandelstam variable for 2-particle scattering. It was then shown in ALrhovac2 that this leads to (2) which we repeat here
+${\rho_{\rm vac}}=\dfrac{3}{4}\dfrac{c^{5}}{\hbar^{3}}\dfrac{m_{\rm z}^{4}}{\mathfrak{g}}\,~~~~~~~~~~(d=3).$ (43)
+For general spatial dimension $d$, the factor of $3/4$ equals $d/(d+1)$ which agrees with (35).
+The derivation of (43) presented in ALrhovac2 from the form-factor bootstrap was new, even in two spacetime dimensions.
+In this derivation of the above formula for $\rho_{\rm vac}$}, since it is based on the properly normalized 2-particle form factor, $m_{\rm z}$ is the physical renormalized mass.
+For integrable theories in $d=1$, the two above definitions of $\rho_{\rm vac}$}, equations (41) and (43), were shown to agree exactly for a wide variety of integrable quantum field theories ALrhovac1 ; ALrhovac2 .
+Various consistency checks on this formula were made, including that it must vanish for supersymmetric theories and also theories with a fractional supersymmetries where the hamiltonian is in the center of the universal enveloping algebra of the conserved charges.
+(See ALrhovac1 and references therein.) The coupling $\mathfrak{g}$ can be positive or negative, and $\rho_{\rm vac}$} can even oscillate in sign for a given model as a function of $\mathfrak{g}$, for instance for the sine-Gordon model where it oscillates between positive and negative, and is zero at the (fractional) supersymmetric points.
+For the sine-Gordon model, there is a special value of the coupling $\mathfrak{g}$ where the fractional supersymmetry is just ${\cal N}=2$ supersymmetry and $\rho_{\rm vac}$} vanishes there as it should ALrhovac1 .
+In 4 spacetime dimensions we did not rigorously prove that equations (41) and (43) are equivalent definitions of $\rho_{\rm vac}$}, thus (2) should still be viewed as well-motivated but still conjectural.
+It’s important to note that the above arguments for the formula (2) do not rely on any kind of spontaneous symmetry breaking.
+For the remainder of the article we simply assume the formula (2).
+In 4 spacetime dimensions, we expect that the “coupling” $\mathfrak{g}$ has a renormalization group flow, based in part on the log’s in (34), and we will assume this in the next section.
+One can motivate this here by considering for instance $\lambda\phi^{4}$ theory where $\mathfrak{g}=\lambda$ to lowest order in perturbation theory.
+Based on (39), one expects that to leading order ${\rho_{\rm vac}}\propto m_{\rm z}^{4}/\lambda$.
+It is known that the marginal coupling $\lambda$ has a non-trivial RG flow Peskin .
+To 1-loop order,
+$\mu\partial_{\mu}\lambda=\dfrac{b}{2\pi}\,\lambda^{2},~~~~~~b=\dfrac{9}{8\pi},$ (44)
+where increasing the energy scale $\mu$ corresponds to a flow to higher energies. In this model, $\lambda$ grows in the flow to high energies, which is to say it is a marginally irrelevant coupling.
+Independent support for the formula (2) comes from Swampland ideas Festina1 ; Festina2 , however this is rather indirect since the arguments are very different as they rely on black holes.
+By studying a charged particle of mass $m$ in the presence of a black hole, it was argued that
+${\rho_{\rm vac}}\leq\dfrac{m^{4}}{2e^{2}}$ (45)
+where here $\mathfrak{g}$ is identified with the quantum electrodynamic (QED) fine structure constant $\alpha=e^{2}/(4\pi\hbar c)$.
+Actual QED will not play a role in this article since we assume the the particle with mass $m_{\rm z}$ is electrically neutral, however, as we will see, it will be instructive to compare couplings and renormalization group parameters with QED, just to check if they are reasonable.
+Our result (2) is stronger than (45) since it is not an inequality, but it is still consistent with (45).
+IV.2 Effective RG flow for Newton’s constant based on the RG flow for $\rho_{\rm vac}$}.
+The formulation of a gravitational Casimir effect in Section II is independent of the detailed properties of $\rho_{\rm vac}$}.
+In this section we use the expression (2) to further develop our dark-universe  model consisting of only vacuum energy.
 For simplicity of subsequent discourse let us rewrite equations (17),(18) explicitly for $d=3$:
@@
 In the expression (2) for $\rho_{\rm vac}$}, by dimensional analysis $\mathfrak{g}$ is a dimensionless coupling constant.
-We assume $m_{\rm z}$ is the physical renormalized mass and ignore its potential RG flow.
-There are good reasons to expect that $\mathfrak{g}$ has an RG flow, based on the comments in ALrhovac1 ; ALrhovac2 ; LeClairUniverse and equation (6) where $\alpha=e^{2}/4\pi\hbar c$ is the fine structure constant with a known RG flow coming from QED.
+Based on the above arguments, $m_{\rm z}$ is the physical renormalized mass of the lightest particle and we thus ignore its potential RG flow.
+Also based on arguments of the last sub-section, we expect that $\mathfrak{g}$ has an RG flow, based on the comments in ALrhovac1 ; ALrhovac2 ; LeClairUniverse and equation (45) where $\alpha=e^{2}/4\pi\hbar c$ is the fine structure constant with a known RG flow coming from QED.
 If we associate $\mathfrak{g}$ with a marginal coupling, $\mathfrak{g}$ could be marginally relevant or irrelevant.
@@
 If $\mathfrak{g}=\alpha=\dfrac{g^{2}}{4\pi\hbar c}$ where $g$ is a gauge coupling with beta-function $\beta_{g}=\dfrac{dg}{d\log\mu}=b\dfrac{g^{3}}{16\pi^{2}}$, then $\beta_{\mathfrak{g}}=b\mathfrak{g}^{2}/2\pi$.
-For pure QED with $N_{f}$ species of Dirac fermions with $g=e$, $\mathfrak{g}_{0}\approx 1/137$ and $b=4N_{f}/3$.
+For pure QED with $N_{f}$ species of Dirac fermions, $\mathfrak{g}_{0}\approx 1/137$ and $b=4N_{f}/3$.
 For $\lambda\phi^{4}$ theory, if $\mathfrak{g}=\lambda$ then $b=9/8\pi$.
@@
 $\mu\partial_{\mu}\mathfrak{g}=\dfrac{b}{2\pi}\mathfrak{g}^{2}.$ (47)
-The above beta function is typically just the 1-loop approximation, however we will argue below, based on cosmological data, that the coupling $\mathfrak{g}$ is presently small so that a 1-loop approximation is justified.
+The above beta function is typically just the 1-loop approximation, however we will argue below based on cosmological data, that the coupling $\mathfrak{g}$ is presently small so that a 1-loop approximation is justified.
+Furthermore, in well-understood QFT’s such as QED, the Landau pole discussed below is not removed by higher order corrections to the beta function.
 We assume $\mathfrak{g}>0$ for a positive $\rho_{\rm vac}$} so that we are dealing with de Sitter space.
@@
 $\dfrac{\mathfrak{g}(\mu)}{\mathfrak{g}_{0}}=\dfrac{1}{1-\widehat{b}\log(\mu/\mu_{0})},~~~~~\widehat{b}\equiv\dfrac{b\,\mathfrak{g}_{0}}{2\pi}.$ (48)
-The RG parameter $\widehat{b}$ is the primary new constant for applications to cosmology.
+The RG parameter $\widehat{b}$ is the primary new constant for applications to cosmology in our framework.
 As we will see, based on cosmological data, we will propose $\widehat{b}>0$, corresponding to a marginally irrelevant coupling, and unless otherwise stated, below $\widehat{b}$ is positive.
@@
 It will play an important role below where we will return to addressing its significance.
+Landau poles are commonplace in elementary particle physics and condensed matter physics, since they are generic to marginally irrelevant perturbations.
+See for instance LandauPole1 ; LandauPole2 ; LandauPole3 and references therein.
+Although QED itself is not relevant to this article, let us mention that it also is not UV complete due to a Landau pole, however in this context it occurs at much higher energies than our proposed $z_{\rm max}$, and is thus commonly viewed as a purely academic issue, and is furthermore complicated by chiral symmetry breaking.
+(See footnote 17).
 From the expression (2) one has
@@
 Thus $\rho_{\rm vac}$} in (46) should be replaced by ${\rho_{\rm vac}}(\mu)$.
-In order to express this equation in terms of ${\rho_{\rm vac}}={\rho_{\rm vac}}(\mu_{0})$ today, it is meaningful to incorporate the RG flow into an induced flow for Newton’s constant:
+In order to express this equation in terms of ${\rho_{\rm vac}}={\rho_{\rm vac}}(\mu_{0})$ today, it is convenient, and meaningful, to incorporate the RG flow into an induced flow for Newton’s constant:
 $\left(\dfrac{\dot{a}}{a}\right)^{2}={\cal H}(t;\mu)^{2}\equiv\dfrac{8\pi{\cal G}(\mu)}{3}{\rho_{\rm vac}}(\mu_{0}),$ (50)
@@
 ${\cal G}(\mu)={\cal G}(\mu_{0})\dfrac{\mathfrak{g}_{0}}{\mathfrak{g}(\mu)}={\cal G}(\mu_{0})\left(1-\widehat{b}\log(\mu/\mu_{0})\right).$ (51)
-It is meaningful to incorporate the RG flow for $\rho_{\rm vac}$} into an induced flow for $G_{N}$ since we view $G_{N}$ as being determined in the far future where vacuum energy dominates and only afterwards incorporate adding matter and energy as we will do below.
+It is meaningful to incorporate the RG flow for $\rho_{\rm vac}$} into an induced flow for $G_{N}$ since we view $G_{N}$ as being determined in the far future where vacuum energy dominates and only afterwards incorporate adding matter and radiation as we will do below.
 If one interprets ${\cal G}(\mu)$ in (51) as an energy scale dependent Newton’s constant, then one sees that for $\widehat{b}>0$, the strength of gravity decreases as one increases the energy scale $\mu$.
@@
 $\dfrac{{\cal H}(t;\mu)}{{\cal H}(t;\mu_{0})}=\sqrt{\dfrac{{\cal G}(\mu)}{{\cal G}(\mu_{0})}}=\sqrt{1-\widehat{b}\log(\mu/\mu_{0})}.$ (52)
-III.2 Tying the energy scale $\mu$ to the temperature in cosmology
+V Exact solution to the scale factor $a(t)$ for the solely dark-universe
 The natural energy scale $\mu$ in cosmology is based on its temperature $T$.
@@
 $T(t)=\dfrac{T_{0}}{a(t)},$ (53)
-since as we will see below the main features will persist with the addition of matter. As is also standard, we express $a(t)$ in terms of the redshift $z$, $a(t)\equiv 1/{(1+z(t))}$, such that
+[…]
 $\mu/\mu_{0}=T/T_{0}=1+z(t).$ (54)
-We thus identify $\mu_{0}=k_{B}T_{0}$, and $z=0$ with $\mu_{0}$. If we identify $G_{N}={\cal G}(\mu_{0})$, then
-${\cal G}(\mu)=G_{N}\left(1-\widehat{b}\log(1+z)\right)=G_{N}\left(1+\widehat{b}\log(a)\right).$ (43)
+This scaling is expected to hold at least post Big Bang Nucleosynthesis, which is roughly a few minutes after the hot big bang.
+Since our model thus far only consists of vacuum energy density $\rho_{\rm vac}$}, we will assume the equation (54) is valid for all times.
+We thus identify $\mu_{0}=k_{B}T_{0}$, and $z=0$ with $\mu_{0}$.
+If we identify $G_{N}={\cal G}(\mu_{0})$, then
+${\cal G}(\mu)={\cal G}(z)=G_{N}\left(1-\widehat{b}\log(1+z)\right)=G_{N}\left(1+\widehat{b}\log(a)\right).$ (55)
 Note that $G_{N}=0$ when $a=a_{\rm min}=e^{-1/\widehat{b}}$.
@@
 Turning to (50), $\dot{a}/a=H$ is no longer constant since $\mu$ is time dependent according to (54). Taking the square-root we wish to solve
-$\left(\dfrac{\dot{a}}{a}\right)=\pm\sqrt{1+\widehat{b}\log a(t)}~\left(\dfrac{8\pi G_{N}\rho_{{\rm vac};0}}{3}\right)^{1/2}$ (44)
-where $\rho_{{\rm vac};0}$ is $\rho_{\rm vac}$} at the scale $\mu_{0}$.
+$\left(\dfrac{\dot{a}}{a}\right)=\pm\sqrt{1+\widehat{b}\log a(t)}~\left(\dfrac{8\pi G_{N}{\rho_{\rm vac}}(\mu_{0})}{3}\right)^{1/2}.$ (56)
 This additional kind of time dependence does not originate directly from Einstein’s equations but rather from the energy scale $\mu$ dependence of Newton’s constant and $\mu$ being tied to the temperature through equation (54).
+By analogy, the classical Maxwell’s equations by themselves cannot account for the well-known energy scale dependence of fine structure constant $\alpha=e^{2}/(4\pi\hbar c)$, which implies it can depend on temperature, and if the temperature is time dependent, such a time dependence is not captured by the time dependence of Maxwell’s equations.
 This leads to a significant and interesting modification of the de Sitter space solution to $a(t)$.
 Let us first chose the positive sign ++ on the RHS of (56) such that $\dot{a}>0$ in the future.
-As we will show, choosing the minus sign will be relevant for the far past.
+As we will show, choosing the minus sign will be relevant for the far past, where the solutions match at $a(t)=a_{\rm min}$.
 The equation (50) can be explicitly integrated, yielding a “inverted gaussian” correction to de Sitter space, which we will simply refer to as gaussian de Sitter space:
@@
 Note that as $\widehat{b}\to 0$, $t_{\rm min}\to$-$\infty$ as expected for de Sitter space.
-In the next section we will approximate $\widehat{b}\approx 0.02$ based on the so-call Hubble tension, which leads to a very large value for $z_{\rm max}$.
+In the next section we will approximate $\widehat{b}\approx 0.02$ based on the so-called Hubble tension, which leads to a very large, but finite value for $z_{\rm max}$ (see eq.
+(66) below.)
 Let us now turn to the far past $t\to$-$\infty$.
-In ordinary de Sitter space, $\lim_{t\to-\infty}a(t)=0$, although it is known this is not a true geometric coordinate singularity.
+In ordinary de Sitter space, $\lim_{t\to-\infty}a(t)=0$, although it is known this is not a true geometric curvature singularity.
 For our gaussian de Sitter space, the solution (57) is formally still valid for $\infty<t<t_{\rm min}$, and thus valid for all times.
@@
 To summarize, the solution (57) is valid for all times $\infty<t<\infty$ where $a(t)>a_{\rm min}$.
-The complete history of such a Universe is that in the far past $t\to$-$\infty$ the scale factor $a(t)$ is infinitely large and starts to compress down to $a_{\rm min}$ which occurs at a time $t_{\rm min}$.
+The complete history of such a model universe, which we referred to as the dark-universe, is that in the far past $t\to$-$\infty$ the scale factor $a(t)$ is infinitely large and starts to compress down to $a_{\rm min}$ which occurs at a time $t_{\rm min}$.
 Beyond this time, $t>t_{\rm min}$, the universe expands until the scale factor $a(t)$ is again infinite at $t=+\infty$.
 This takes an infinite amount of time, and avoids any geometric singularities at $a=0$.
-There is no “Big Bang” that occurs at time $t_{\rm min}$, rather at this time the Universe smoothly transitions from a compression to an expansion, and $t_{\rm min}$ is just the time when the Universe is hottest.
-The Big Bang is now better described as a Swing.
-An picturesque analogy is the harmonic pendulum with the bottom of the swing corresponding to $a_{\rm min}$, and at the present time the Universe is on the up swing.
-Based on (49), a consistent possibility is that this process then repeats itself, such that Universe is in a sense oscillating about its vacuum like a pendulum harmonic oscillator.
+There is no “Big Bang” corresponding to scale factor $a(t)=0$.
+Rather, at time $t_{\rm min}$ the universe smoothly transitions from a compression to an expansion, and $t_{\rm min}$ is just the time when the universe is hottest.
+The Big Bang is now better described as a Big Swing.
+Henceforth, by “big bang” we refer to the time $t_{\rm min}$ which is a very hot big bang and not associated with any curvature singularities at $a=0$.
+A picturesque analogy is the harmonic pendulum with the bottom of the swing corresponding to $a_{\rm min}$, and at the present time the universe is on the up swing.
+When we add quantum matter and radiation in the next section, these are viewed as excitations above the vacuum, and as the pendulum is on the downswing it accumulates kinetic energy which can excite these states.
+Based on (61), a consistent possibility is that this process then repeats itself, such that this universe is in a sense oscillating about its vacuum like a pendulum harmonic oscillator about its ground state.
 At the top of the swing the universe is at its largest.
-If this is the case, then the age of this gaussian de Sitter Universe is even more infinite.
-IV Adding matter: the $\Lambda{\rm CDM}$ model re-examined and the Hubble Tension
-Thus far we have only considered vacuum energy density $\rho_{\rm vac}$}.
+If this is the case, then the age of this gaussian de Sitter universe is even more eternal.
+VI Adding matter: the $\Lambda{\rm CDM}$ model re-examined and the Hubble Tension
+Thus far we have only considered vacuum energy density $\rho_{\rm vac}$} which led to the exact solution of the dark-universe presented above.
 We gave a heuristic derivation of Einstein’s equations in this context from the gravitational Casimir effect we formulated above.
 The Universe also contains radiation and matter, which we view as excitations above the vacuum, which must be included in $T_{\mu\nu}$.
-Local conservation requires vanishing of the covariant divergence $\nabla^{\mu}T_{\mu\nu}=0$, which is automatic due to the Bianchi identities.
-Thus, for instance, Newton’s universal law for the gravitational force between two masses is subsumed as a consequence.
-At this point one must deal with the usual Einstein’s equations except with an energy dependent Newton’s constant ${\cal G}(\mu)$:
+The standard Friedmann equations have three sources, as in (63) below, but without the $(1+\widehat{b}\log a)$ factor.
+The question naturally arises: should the renormalization properties of $\rho_{\rm vac}$} be kept to the $\Omega_{\Lambda}$ only, i.e.
+the $\rho_{\rm vac}$} term, or should it be incorporated also in the $\Omega_{\rm m}$ and $\Omega_{\rm rad}$ terms?
+We take the following point of view, although it perhaps requires more scrutiny, or at least a more complete argument.
+The dark-universe  is considered a skeleton of our Universe with pure vacuum energy $\rho_{\rm vac}$} and this led to the emergence of $G_{N}$ from $\rho_{\rm vac}$}.
+Matter and radiation are excitations over the vacuum, however we continue to assume that $G_{N}$ is determined by vacuum energy since we argued that the latter is the origin of Gravity itself.
+In our analysis of the dark-universe, the renormalization group flow coming from $\rho_{\rm vac}$} was absorbed into an energy scale $\mu$ dependent Newton’s constant $G_{N}$.
+If the energy scale $\mu$ is fixed, then we add matter and radiation by demanding local energy-momentum conservation, which requires vanishing of the covariant divergence $\nabla^{\mu}T_{\mu\nu}=0$, which is automatic due to the Bianchi identities.
+Thus, for instance, Newton’s universal law for the gravitational force between two masses at low energies is subsumed as a consequence.
+At this point one must deal with the usual Einstein’s equations except with an energy dependent Newton’s constant ${\cal G}(\mu)$.
+Based on this we propose that:
 $\left(\dfrac{\dot{a}}{a}\right)^{2}=\dfrac{8\pi}{3}{\cal G}(\mu)\,\rho_{\rm total},$ (62)
-with ${\cal G}(\mu)$ defined in (43). For the moment, let us not tie the energy scale to $\mu$ to $z(t)$ as we did in (42). It is conventional to express (50) as follows
+with ${\cal G}(\mu)$ defined in (55). It is conventional to express (62) as follows
 $H^{2}\equiv\left(\dfrac{\dot{a}}{a}\right)^{2}=H_{0}^{2}\left(1+\widehat{b}\log a\right)\left(\dfrac{\Omega_{\rm rad}}{a^{4}}+\dfrac{\Omega_{\rm m}}{a^{3}}+\Omega_{\Lambda}\right),$ (63)
 where by definition $a(t_{0})=1$ at the present time $t_{0}$, $H(t_{0})=H_{0}$, and $\Omega_{\Lambda}$ is the $\rho_{\rm vac}$} contribution.
+As for the dark-universe, the additional time dependence due to the factor $(1+\widehat{b}\log a(t))$ does not follow directly from the Einstein equations, but rather from the energy scale $\mu$ dependence of Newton’s constant and $\mu$ being tied to the temperature through equation (54).
+The Bianchi identities are only valid locally where $\mu$ is a fixed and thus Newton’s constant ${\cal G}(\mu)$ is time independent.
 The zero curvature ($k=0)$ motivated in Section II implies $\Omega_{\rm rad}+\Omega_{\rm m}+\Omega_{\Lambda}=1$.
+The above prescription (63) for dealing with the issue raised preserves the minimal scale factor $a_{\rm min}$ as we will show below.
+It is also a key aspect of our proposal for dealing with the Hubble tension.
 In the $\Lambda{\rm CDM}$ model one ignores radiation, i.e.
@@
 The first comes from “local” measurements based on Type Ia supernovae, where typical redshifts are relatively low, in the range z=0.020.15z=0.02–0.15 (SHOES SHOES .).
-The other determination of $H_{0}$ is based on Cosmic Microwave Background data at much higher redshift $z=1100$.
+The other determination of $H_{0}$ is based on CMB data at much higher redshift $z=1100$.
 (Planck Planck ).
-The two contradictory values reported are
+The two contradictory values currently reported are
 $H_{0;{\rm CMB}}=67.4\pm 0.5~{\rm km/s/Mpc},~~~~~H_{0;{\rm SN}}=73.0\pm 1.0~{\rm km/s/Mpc},$ (64)
-which differ significantly enough for some prominent researchers to question the $\Lambda{\rm CDM}$ model, with suggestions that this could signify beyond standard model physics HubbleTensionRef .
-In the latter it was suggested that this could be due to a variable Dark Energy component, but no specific theoretical model was proposed.
+which differ significantly enough for some prominent researchers to question the $\Lambda{\rm CDM}$ model, with suggestions that this could signify beyond standard model physics HubbleTensionRef ; Riess2 .
+In the latter it was suggested that this could be due to a variable Dark Energy component, but no specific theoretical model was advocated.
 Our interpretation of the discrepancy in these two values for $H_{0}$ in (64) is the following.
@@
 We thus propose based on (52)
-$\dfrac{H_{0;{\rm CMB}}}{H_{0;{\rm SN}}}\approx\sqrt{1-\widehat{b}\log(1+z)},~~~~{\rm for}~z=z_{\rm CMB}=1100,$ (53)
+$\dfrac{H_{0;{\rm CMB}}}{H_{0;{\rm SN}}}\approx\sqrt{1-\widehat{b}\log(1+z)},~~~~{\rm for}~z=z_{\rm CMB}=1100~~\Longrightarrow~~\widehat{b}\approx 0.02$ (65)
 where we have set $z\approx 0$ for the supernovae measurements. Based on the measured values in (64), the above leads to $\widehat{b}\approx 0.02$.
@@
 This is a robust conclusion, regardless of whether $\rho_{\rm rad}$ is incorporated as the dominant form of energy in the very early Universe.
-The above value for $\widehat{b}$, leads to a very high value for the maximal red shift: [^17] This is not an unreasonable value for $\widehat{b}$, in that $\mathfrak{g}_{0}=2\pi\widehat{b}/b=0.12/b$, and typically $b={\cal O}(1)$.
+The above value for $\widehat{b}$, leads to a very high value for the maximal red shift and temperature: [^19] This is not an unreasonable value for $\widehat{b}$, in that $\mathfrak{g}_{0}=2\pi\widehat{b}/b=0.12/b$, and typically $b={\cal O}(1)$.
 For comparison, for QED $b=4N_{f}/3$ for $N_{f}$ species of Dirac fermions, and if $\mathfrak{g}_{0}=1/137$ then $\widehat{b}=.002N_{f}$ which is about 10 times smaller than in (66) and $z_{\rm max}\approx 10^{500}$ is far beyond the Planck scale.
-$\widehat{b}\approx 0.02~~~\Longrightarrow~~a_{\rm min}\approx 2\times 10^{-22},~~~~z_{\rm max}\approx 5\times 10^{21}.$ (54)
+$\widehat{b}\approx 0.02~~~\Longrightarrow~~a_{\rm min}\approx 2\times 10^{-22},~~~~z_{\rm max}\approx 5\times 10^{21}~~~\Longrightarrow~~~T_{\rm max}\approx 10^{22}\,K.$ (66)
 This is far above what has been directly probed by the CMB, the latter being $z_{\rm CMB}=1100$.
@@
 For instance it doesn’t significantly alter what was previously considered as the age of the Universe, as we will show below.
-To further justify the last statement, let us turn to the $\Lambda{\rm CDM}$ model where radiation $\Omega_{\rm rad}=0$ as a good approximation,.
-In order to probe potential singularities at early times where the matter dominates due to the $1/a^{3}$ in (50), let us neglect the $\Omega_{\Lambda}$ term as an approximation.
+To further justify the last statement, let us turn to the $\Lambda{\rm CDM}$ model where radiation $\Omega_{\rm rad}=0$ is a good approximation.
+The scale factor $a(t)$ can be solved numerically starting from (63).
+In order to probe potential singularities at early times where the matter dominates due to the $1/a^{3}$ in (62), let us neglect the $\Omega_{\Lambda}$ term as an approximation, since the resulting equation can be solved analytically and leads to more transparent conclusions.
 Then one wishes to integrate the equation
@@
 There exists a time $t_{\rm min}$ where aa=0: [^20] In the cosmology literature the convention is to shift $t\to t+t_{\rm min}$ such that $a=0$ at the shifted time $t=0$.
-The difference $t_{0}-t_{\rm min}$ essentially represents the age of the Universe if one assume the Universe came into existence at time $t_{\rm min}$, which is questioned in this article.
+The difference $t_{0}-t_{\rm min}$ essentially represents the age of the Universe if one assumes the Universe came into existence at time $t_{\rm min}$, which is questioned in this article.
 $a(t_{\rm min})=0~~{\rm for}~~t_{\rm min}=t_{0}-\dfrac{2}{3H_{0}\sqrt{\Omega_{m}}}.$ (69)
@@
 Thus due to the square-root branch cut $\sqrt{1+\widehat{b}\log a}$ inside $\rm erfi$}, there is a minimal value of $a(t)$ again given by (58).
+It is not difficult to check that the same conclusion is reached in a radiation dominated era where the $1/a^{3}$ term is replaced by $\Omega_{\rm rad}/a^{4}$.
 This minimal value of $a(t)$ occurs at a time $t_{\rm min}$ where $a(t_{\rm min})=a_{\rm min}$, which is determined by when the LHS of the above equation is zero since ${\rm erfi}(0)=0$:
 $a(t_{\rm min})=a_{\rm min}=e^{-1/\widehat{b}}$ (71)
-where
+where $t_{\rm min}$ is determined by the equation
 $t_{0}-t_{\rm min}=\dfrac{1}{H_{0}}e^{-3/2\widehat{b}}\sqrt{\dfrac{2\pi}{3\widehat{b}\,\Omega_{m}}}{\rm erfi}\sqrt{\dfrac{3}{2\widehat{b}}}=\dfrac{2}{3H_{0}\sqrt{\Omega_{m}}}\left(1+\tfrac{1}{3}\,\widehat{b}+\tfrac{1}{3}\,\widehat{b}^{2}+\tfrac{5}{9}\,\widehat{b}^{3}+\ldots\right).$ (72)
@@
 Using ${\rm erfi}(-x)=-{\rm erfi}(x)$, one can show that equation (61), which equates $a(t)$ with $a(-t+2t_{\rm min})$, remains valid with $t_{\rm min}$ given in (72), as does (60).
-Based on the above analysis of our scenario, let us comment on the possible implications for the very early Universe.
-Henceforth, by “early Universe” we mean around the time $t_{\rm min}$, where the expansion phase begins.
+Based on the above analysis, let us comment on the possible implications for the very early universe in our model.
+Henceforth, by “early universe” we mean around the time $t_{\rm min}$, where the universe is hottest and the expansion phase begins.
 Recall in our model, time $t$ actually extends to $t=$-$\infty$, and this is the earliest universe.
-The salient features of the evolution of $a(t)$ from $t=$-$\infty$ to $t=+\infty$ is essentially the same as for the pure vacuum energy case, described in the last paragraph of the last section.
+The salient features of the evolution of $a(t)$ from $t=$-$\infty$ to $t=+\infty$ is essentially the same as for the pure vacuum energy case, i.e.
+the dark-universe  described in the last paragraph of the last section, and this justifies viewing the dark-universe  as a skeleton of our universe based only on the Vacuum.
 Namely at $t=$-$\infty$, $a(t)=\infty$ and starts to decrease until it reaches $a_{\rm min}$ at time $t_{\rm min}$, then begins to increase back to $\infty$ as $t\to+\infty$.
@@
 There is no singular Big Bang at any time, rather at time $t_{\rm min}$ the universe is hottest since the redshift $z$ is at its maximum.
-The quantity $t_{0}-t_{\rm min}$ does not represent the age of the Universe, but rather the time elapsed since the scale factor was at its minimum.
-The age of the Universe is actually infinite, which we find appealing, since the Universe does not originate from an incomprehensible singularity at $a=0$.
-As described above for the case of only Vacuum energy, the time evolution of the Universe is better described as a back and forth swing of a pendulum, where the bottom of the swing occurs at $a(t_{\rm min})=a_{\rm min}$, and we are currently on an upswing.
-The details of this evolution are very sensitive to the value of $\widehat{b}$ due to the exponentials in (54).
-For purposes of illustration, let us assume $\widehat{b}=0.02$ as inferred from the Hubble tension.
+The quantity $t_{0}-t_{\rm min}\approx 13.7$ billion years, does not represent the age of the universe, but rather the time elapsed since the scale factor was at its minimum.
+The age of this model universe is actually infinite, which we find appealing, since the universe does not originate from an incomprehensible singularity at $a=0$.
+As described above for the dark-universe, the time evolution of the universe is better described as a back and forth swing of a pendulum, where the bottom of the swing occurs at $a(t_{\rm min})=a_{\rm min}$, and we are currently on an upswing.
+All the energy of the hot universe at time $t_{\rm min}$ was accumulated during the downswing, thus such a universe was not created out of nothing, since the Vacuum is not nothing; it is something with a ground state energy density that is a source of energy, if not the original source of all energy in the Universe.
+The details of this time evolution of $a(t)$ are very sensitive to the value of $\widehat{b}$ due to the exponentials in (66).
+For purposes of illustration, let us assume $\widehat{b}=0.02$ as inferred above from the Hubble tension.
 First, one sees from (72) that for small $\widehat{b}$ this amounts to a small corrections to $t_{0}-t_{\rm min}$ which is normally associated to the age of the Universe.
-If our estimate of $\widehat{b}$ in (54) is approximately correct, then this value for $z_{\rm max}$ is deeply into the radiation dominated era, since the value of $z=z_{\rm eq}$ where matter and radiation are equivalent in density is only about $z_{\rm eq}\approx 3400$.
-At times much earlier than during the $\Lambda{\rm CDM}$ era, equation (42) is not necessarily valid, however if we simply assume it extends to the Planck scale with $T_{0}=2.7$K, then $z_{\rm planck}\approx 5\times 10^{31}$, which is much higher than $z_{\rm max}$ by 10 orders of magnitude.
-This indicates that the Planck scale plays no role in determining $z_{\rm max}$, which is to say quantum gravity does not play a role.
-For our perspective on gravity, earlier times $t<t_{\rm min}$ where $a(t)<a_{\rm min}$ do not exist in our model.
-In the standard cosmology this earlier era is thought to be described by inflation.
+If our estimate of $\widehat{b}$ in (66) is approximately correct, then this value for $z_{\rm max}\approx 5\times 10^{21}$ is deeply into the radiation dominated era, since the value of $z=z_{\rm eq}$ where matter and radiation are equivalent in density is only about $z_{\rm eq}\approx 3400$.
+At times much earlier than during the $\Lambda{\rm CDM}$ era, equation (54) may not be exact at all times, since new degrees of freedom can be excited from the Vacuum and go through various phase transitions that modify (54) in relatively short in time epochs.
+However if we simply assume it extends to the Planck scale with $T_{0}=2.7$K, then $z_{\rm planck}\approx 5\times 10^{31}$, which is much higher than $z_{\rm max}$ by 10 orders of magnitude.
+This shows that the Planck scale $\ell_{p}$ plays no role in determining $z_{\rm max}$.
+However a smaller value for $\widehat{b}$ could easily raise $z_{\rm max}$ up to the Planck scale, where quantum gravity effects could modify our findings.
+From the perspective on Gravity presented thus far, earlier times $t<t_{\rm min}$ where $a(t)<a_{\rm min}$ do not exist in our model universe.
+In the standard cosmology, this earlier era is thought to be described by inflation.
 Inflation is postulated as a way to get from the singularity at $a=0$ to a hot big bang, however it is not yet universally accepted as being a complete theory.
+It requires invoking new fields such as the inflaton and fine-tuning their parameters for a “slow-roll”, incorporating re-heating, etc.
 Very large numbers are naturally involved since the singularity at $a=0$ formally corresponds to $z=\infty$.
-Typical models require inflation to last for 50-60 e-foldings to inflate the Universe to the proper initial conditions from $a\approx 0$, then transition to a re-heating scale to bring up the temperature to that of a hot Big Bang, and this rapid expansion makes the Universe flat and homogeneous.
-The range of redshift $z$ over which inflation hypothetically occurs is strongly model dependent, especially on what the re-heating temperature is, and the duration of inflation, marking its beginning and end, can range from $10^{22}<z_{\rm inflation}<10^{53}$, where the upper limit can even be significantly larger than $z_{\rm Planck}$.
-[^19] During hypothetical inflation, the redshift $z$ is not given by the formula (42), but based on the definition of the scale factor $a=1/(1+z)$.
+Typical models require inflation to last for 50-60 e-foldings to inflate the Universe to the proper initial conditions from $a=0$, then transition to a re-heating scale to bring up the temperature to that of a hot big bang, and this rapid expansion makes the Universe flat and homogeneous.
+The range of redshift $z$ over which inflation hypothetically occurs is strongly model dependent, especially on what the re-heating temperature is.
+The duration of inflation, marking its beginning and end, can range from $10^{22}<z_{\rm inflation}<10^{53}$ depending on the model, where the upper limit can even be significantly larger than $z_{\rm planck}$.
 It is noteworthy that at the lower end of this range based on lower reheating temperatures, namely where inflation ends at $z_{\rm inflation}=10^{22}$, is close to our $z_{\rm max}=5\times 10^{21}$.
-Since this inflationary epoch does not actually exist in our model, one should question whether inflation is necessary in the standard model of cosmology.
-At $t=t_{\rm min}$ the Universe is already expanded since $a(t)>a_{\rm min}$, and the Universe is already flat with $k=0$.
+Since this inflationary epoch does not actually exist in our model due to the minimal scale factor $a_{\rm min}$, one should question whether inflation is really necessary.
+At $t=t_{\rm min}$ our model universe is already expanded since $a(t)>a_{\rm min}$, it is very hot, and is already flat with $k=0$.
 In Section II we explained why the spatial curvature $k$ is naturally zero since we view Gravity as arising from vacuum energy in flat Minkowski space.
 The value of $a_{\rm min}=e^{-1/\widehat{b}}$ perhaps plays a role equivalent to hypothetical inflation models since it represents roughly 50 e-foldings since $e^{1/\widehat{b}}\approx e^{50}$ for our estimate of $\widehat{b}\approx 0.02$.
-V Closing Remarks
-Having already summarized the unconventional perspective on Gravity developed in this article and its implications in the Abstract and Introduction, we close with some suggestions for further investigations and discuss some open questions.
-There are currently many open avenues for testing (and falsifying) our main proposals.
-$\bullet$   Above we considered implications for the cosmology based on the $\Lambda{\rm CDM}$ model.
-The energy scales involved are relatively low since $\mu_{0}=k_{B}T_{0}=2.35\times 10^{-4}\,{\rm eV}$ for $T_{0}=2.7$ K.
-If our proposed energy dependent Newton’s constant ${\cal G}(\mu)$ in equation (43) is correct, then it is in principle possible to confirm it in some bench-top Cavendish like experiments as a function of temperature.
-Suppose such experiments are carried out in a small range of temperature $\Delta T$ about a fixed temperature $\widetilde{T}$.
-From (39)
-${\cal G}(T)={\cal G}(\widetilde{T})\left(1-\widehat{b}\log(T/\widetilde{T})\right)\approx{\cal G}(\widetilde{T})\left(1-\dfrac{\widehat{b}}{\widetilde{T}}\,\Delta T\right).$ (61)
-As discussed above, the effective Newton’s constant decreases with increased temperature.
-If $\widetilde{T}=300$ K is room temperature, then $\widehat{b}/\widetilde{T}\approx 6\times 10^{-5}$ per degree change $\Delta T$ if $\widehat{b}=0.02$, where this value of $\widehat{b}$ was inferred from the Hubble tension in (53).
-There are relatively few experiments along these lines, going back as far as 1923 Shaw ; Dmitriev ; Liangzao , wherein weight is measured as a function of temperature for various metals around room temperature.
-These experiments indeed consistently measure a decrease in the gravitational force as the temperature is increased, as our model predicts.
-Based on the ideas in this article, this decrease should be independent of the material, in this case a metal, since the Equivalence Principle is built into the formalism, whereas experiments find slightly different relative changes to $G_{N}$ depending on the metal.
-This could be due to experimental uncertainties as a result of buoyancy, thermal expansion, and other factors which are difficult to take into account.
-In any case a linear fit of this experimental data in $\Delta T$ according to (61) yield variations of about $10^{-5}$ per degree, which is roughly consistent with our estimate of $\widehat{b}/\widetilde{T}$ for $\widehat{b}\approx 0.02$ and $\widetilde{T}=300$K.
-[^20] The article Liangzao does not attempt a fit to their data.
-We made a crude fit of their tabular results to the functional form (61) and found $\widehat{b}/\widetilde{T}$ to be $0.8\times 10^{-5}$ for gold and $1.4\times 10^{-5}$ for nickel, however a closer inspection and interpretation of this data beyond our crude fit is warranted.
-At lower reference temperatures $\widetilde{T}$, the effect is larger and this motivates carrying out such Cavendish type experiments at lower temperatures.
+On the other hand, one must be careful not to discard the successes of inflationary models, in particular their hallmark prediction of quantum density fluctuations that seed future galaxies and has been observed in the CMB.
+In our model universe, such quantum fluctuations could arise from the field corresponding to the particle with mass $m_{\rm z}$ itself, which should be interpreted as doing the job of the inflaton.
+Clearly this requires further study beyond the original scope of this article.
+VII Consistency with various types of constraints and experiments
+In this section we compare the primary aspects of our model to known observations and experiments of various types.
+We will base our comparison based on our estimate $\widehat{b}\approx 0.02$, however conclusions are rather sensitive to the value of $\widehat{b}$ and could easily be accommodated with a smaller $\widehat{b}$.
+VII.1 $H_{0}(z)$ trend for low $z$ Supernovae
+Since (65) is only based on the two data points $z=0$ and $1100$, more compelling support for our proposal would be a fit for a range of redshift $z$ that would confirm the trend proposed above:
+$\dfrac{H_{0}(z)}{H_{0}(z=0)}=\sqrt{1-\widehat{b}\log(1+z)}=1-\dfrac{\widehat{b}}{2}\,z+\dfrac{\widehat{b}(2-\widehat{b})}{8}\,z^{2}+{\cal O}(z^{3}).$ (73)
+This is possible with supernovae since they exist in a range of redshifts.
+Remarkably this trend has been very recently observed for a large data sample of supernovae in the range of still relatively low $z$, $0.001<z<2.3$, based on data analyzed at the National Astronomical Observatory of Japan Dainotti2021 ; Dainotti2022 ; Dainotti2024 ; Dainotti2025 .
+In the latter work, a fit to the functional form
+$\dfrac{H_{0}(z)}{H_{0}(z=0)}\approx\dfrac{1}{(1+z)^{\alpha}}=1-\alpha\,z+\dfrac{\alpha(1+\alpha)}{2}\,z^{2}+{\cal O}(z^{3})$ (74)
+was considered, motivated by some versions of $f(R)$ modified gravity and other theoretical models which typically invoke additional fields, for instance Brans-Dicke scalars.
+(See Montani ; Montani2 ; Dainotti2021 ; Dainotti2022 ; Dainotti2024 ; Dainotti2025 and references therein.) Although the two formulas (73) and (74) look rather different, to lowest order in small $z$, they agree precisely with $\alpha=\widehat{b}/2$.
+In Dainotti2021 ; Dainotti2022 ; Dainotti2024 ; Dainotti2025 values of $\alpha\approx 0.01$ were reported, which agrees with our estimate of $\widehat{b}\approx 0.02$ inferred from the single CMB data point $z=1100$ in (65).
+This strongly suggests that the trend in $H_{0}(z)$ seen in supernovae extends all the way to $z=1100$, and this provides the strongest support for our model thus far.
+[^21] We are currently working with Maria Dainotti’s group at NAOJ to perform a detailed statistical analysis in order to try and distinguish between the two formulas (73) and (74).
+Preliminary results appear to validate our estimate of $\widehat{b}\approx 0.02$ ALMaria .
+It clearly would be very interesting to extend this range of analysis to supernovae with higher $z$ or to gamma ray bursts.
+VII.2 Known bounds on the time-variation of Newton’s constant
+If the energy scale varying $G_{N}$ proposed in this article is correct, it is testable by other kinds of measurements.
+For cosmology, our formula (51) for the scale dependent Newton’s constant ${\cal G}(\mu)$ implies that the effective Newton’s constant is time dependent if one ties $\mu$ to the time-dependent scale factor as in (54).
+Unfortunately, presently there only exists experimental bounds on the time variation of $G_{N}$, usually reported as $\dot{G}_{N}/G_{N}$ today, rather than non-null results.
+Nevertheless, let us check consistency with some known bounds.
+Based on (55) one has the simple formula
+$\dfrac{\dot{G}_{N}}{G_{N}}(t)=H(t)\cdot\widehat{b}$, (75)
+where as above $H=\dot{a}/a$. At the present time $t_{0}$,
+$\dfrac{\dot{G}_{N}}{G_{N}}=H_{0}\cdot\widehat{b}=6.7\times 10^{-11}\cdot\widehat{b}~/{\rm yr}~~~~({\rm today}),$ (76)
+where we have used $H_{0;{\rm CMB}}$. Henceforth, $\dot{G}_{N}/G_{N}$ refers to the present time.
+Pulsar timing and gravitational waves.
+The strongest experimental constraints come from pulsar timing Taylor based on pulsar observations spanning 30 years Taylor .
+Over this rather narrow window, one can constrain $\dot{G}_{N}/G_{N}<2\times 10^{-12}/{\rm yr}$.
+These pulsars exist at relatively low redshift $z$.
+Thus based on equation (76) with $\widehat{b}=0.02$, $\dot{G}_{N}/G_{N}\approx 1.4\times 10^{-12}\,/{\rm yr}$ which is very close to the bound based on pulsars.
+LIGO provides weaker constraints $10^{-9}/{\rm yr}$ LIGO .
+Big Bang Nucleosynthesis    The above results indicate that our model is not yet ruled out up to the CMB scale at $z\approx 1100$.
+This is still a relatively low energy scale corresponding to a temperature 30003000K.
+As stated in the Introduction, we assume that (73) extends to higher $z$, and this necessarily involves indirect inferences and consequently will be less conclusive.
+The next highest energy scale that can provide constraints is Big Bang Nucleosynthesis (BBN) at $z\approx 10^{9}$, corresponding to an energy of about 11  MeV and a temperature of about $10^{9}$K, whose consequences can be inferred from abundances of light nuclei such as Li.
+Based on the formula (55), at this relatively high $z$, $G_{N}$ is reduced by about 40% if $\widehat{b}=0.02$.
+The constraints on the time variation of $G_{N}$ for BBN are typically formulated as $\Delta G_{N}/G_{N}$, where $\Delta G_{N}$ is a bound on how much $G_{N}$ can differ at the time of BBN.
+It’s important that the reported bounds assume slow or no variation after BBN, and $\dot{G}_{N}/G_{N}$ is based solely on a linear fit between the time of BBN and today, and this should be kept in mind in drawing conclusions.
+Incorporating the full time evolution based on (55) could significantly alter change these bounds.
+Nevertheless, let us proceed with some simple checks.
+The earliest article is Krauss concludes that to $1\sigma$ confidence level, $0.85<G_{\rm BBN}/G_{N}<1.21$, which allows a 20% change of $G_{N}$ in either direction in the BBN epoch, which is borderline consistent with our 40% decrease; as we stated a small change in $\widehat{b}$ could potentially accommodate this.
+To compare with the prediction (76), one should convert this to a time variation over the intervening time, and the same article reports $\dot{G}_{N}/G_{N}<4\times 10^{-13}\,{\rm yr}^{-1}$ which potentially conflicts with $1.4\times 10^{-12}\,/{\rm yr}$ based on $H_{0}$ and $\widehat{b}=0.02$.
+More recently, tighter bounds were reported by inferring the consequences of BBN on the CMB Alvey .
+With $2\sigma$ confidence level, this article reports $0.94<G_{\rm BBN}/G_{N}<1.05$, which translated to a time evolution one obtains $\dot{G}_{N}/G_{N}<4.5\times 10^{-12}\,{\rm yr}^{-1}$ Alvey .
+The latter is consistent with our $1.4\times 10^{-12}\,/{\rm yr}$ estimate above.
+We conclude that more work is needed here to determine whether our model is consistent with models of BBN, in particular the effect of the complete time evolution which incorporates the time varying $G_{N}$ implicit in (63).
+VII.3 Before Big Bang Nucleosynthesis?
+At much higher $z$, experimental probes are more severely limited and indirect.
+Nevertheless some observations are worthwhile to point out.
+As previously discussed, if the Hubble tension trend extends to this very early universe, then based on our estimate of $\widehat{b}\approx 0.02$, there is a minimal scale factor $a_{\rm min}$ and a corresponding $z_{\rm max}\approx 5\cdot 10^{21}$, which corresponds to a hottest temperature $T_{\rm max}\approx 10^{22}$K, which we associated with a hot big bang.
+First of all, at least this $T_{\rm max}$ is above the electro-weak scale of about $160$ GeV which is a temperature of about $10^{15}$K, otherwise the electro-weak transition would have never occurred.
+The temperature $T_{\rm max}$ is roughly the energy scale of some Grand Unified Theories, depending on the model.
+In the last section we explained how $z_{\rm max}$ is roughly the scale for low-scale inflation models, however our model does not exist for scale factors $0<a(t)<a_{\rm min}$, which is normally considered the inflationary era.
+This led us to suggest that current models of inflation may not be necessary, since at $z_{\rm max}$ the Universe is already expanded, hot, and flat.
+Furthermore, before the time when $a(t)=a_{\rm min}$, our solution to $a(t)$ is valid due to (61), namely $a(t)=a(-t+2t_{\rm min})$.
+A further constraint on our model is that it should not spoil other positive predictions of inflation, in particular how quantum density fluctuations of the hypothetical inflaton field can explain the primordial density fluctuations which are thought to provide the seeds for large scale galaxy formation and probed by the CMB.
+Quantum fluctuations of the field for lightest particle of mass $m_{\rm z}$, upon which our formula for $\rho_{\rm vac}$} in (2) is based upon, could potentially play such a role, but such considerations require further investigation.
+VII.4 Bench-top experiments on the temperature dependence of Newton’s constant
+Above we considered implications for cosmology based on the $\Lambda{\rm CDM}$ model and beyond.
+Apart from the trend recently observed for supernovae discussed above in subsection A, unfortunately we could only compare with bounds rather than definitive non-null experimental signatures.
+The formula (55) implies that Newton’s constant actually depends on temperature.
+An analogy can be made with QED where one must take into account the RG flow of the fine structure constant to make predictions at higher energies.
+For the $\Lambda{\rm CDM}$ model, the energy scales involved are relatively low since $\mu_{0}=k_{B}T_{0}=2.35\times 10^{-4}\,{\rm eV}$ for $T_{0}=2.7$ K.
+For cosmology, the temperature $T$ as described above is the temperature of the entire Universe, thus it is not clear if such a temperature dependent Newton’s constant is also valid locally in space or time.
+If it is, and our proposed energy dependent Newton’s constant ${\cal G}(\mu)$ in equation (55) is correct, then it is in principle possible to confirm it in some bench-top Cavendish like experiments as a function of temperature.
+If this could be confirmed, it would be truly remarkable that the new fundamental parameter $\widehat{b}$ introduced in this article based on the Hubble tension could be measured in a small bench-top experiment.
+Ideally one wishes to measure the gravitational force between two masses as a function of their equal temperatures, however we found no such such experiments in the literature.
+We did find 3 not so well-known experiments wherein the weight of a sample is measured as a function of temperature Shaw ; Dmitriev ; Liangzao , the earliest going as far back as 1923, wherein a positive result is reported rather than a bound.
+Since only the temperature of the sample is varied, and not that of the Earth, it’s again not clear if our proposal applies directly as stated, however let us proceed.
+Suppose such experiments are carried out in a small range of temperature $\Delta T$ about a fixed temperature $\widetilde{T}$, i.e. $T=\widetilde{T}+\Delta T$. From (51)
+${\cal G}(T)={\cal G}(\widetilde{T})\left(1-\widehat{b}\log(T/\widetilde{T})\right)={\cal G}(\widetilde{T})\left(1-\dfrac{\widehat{b}}{\widetilde{T}}\,\Delta T+\dfrac{\widehat{b}}{2\widetilde{T}^{2}}\,(\Delta T)^{2}+\ldots\right),$ (77)
+[…]
+VIII Closing Remarks
+Having already summarized the unconventional perspective on Gravity developed in this article and its implications in the Introduction, we close with some suggestions for further investigations and discuss some open questions.
+There are currently many open avenues for testing (and falsifying) our main proposals, in addition to those considered in the last section.
 $\bullet$   The induced RG flow for Newton’s constant that we proposed as a consequence of the RG flow of the coupling $\mathfrak{g}$ in $\rho_{\rm vac}$} in (2) led to an avoidance of a geometric curvature singularity at a scale factor $a=0$.
@@
 This could potentially lead to a resolution of the singularity at the origin of the Schwarzschild black hole solution, in a way analogous to the minimal scale factor $a_{\rm min}$ above.
-$\bullet$   If the varying $G_{N}$ proposed in this article is correct, it is testable by other kinds of measurements.
-Let us speculate on some possible directions or topics for further study.
-We already mentioned above that in principle this varying $G_{N}$ can be measured in bench-top Cavendish experiments as a function of temperature, and that the very few existing experiments carried out at room temperature are roughly consistent with the parameters we inferred from the Hubble tension.
-For cosmology, our formula (39) for the scale dependent Newton’s constant ${\cal G}(\mu)$ implies that the effective Newton’s constant is effectively time dependent if one ties $\mu$ to the time-dependent scale factor as in (42).
-There are known experimental bounds on the time variation of $G_{N}$, usually reported as $\dot{G}_{N}/G_{N}$.
-Based on (43) one has
-$\dfrac{\dot{G}_{N}}{G_{N}}=H\,\widehat{b}$, (62)
-where as above $H=\dot{a}/a$.
-The strongest experimental constraints come from pulsar timing Taylor which constrains $\dot{G}_{N}/G_{N}<2\times 10^{-12}/{\rm yr}$.
-LIGO provides weaker constraints $10^{-9}/{\rm yr}$ LIGO .
-These pulsars exist at relatively low redshift $z$.
-Thus taking $H=H_{0}$ in the above equation (62), one finds $\dot{G}_{N}/G_{N}\approx 7\times 10^{-11}\times\widehat{b}\,/{\rm yr}$.
-For $\widehat{b}\approx 0.02$, as we proposed based on Hubble tension data, $\dot{G}_{N}/G_{N}\approx 1.4\times 10^{-12}\,/{\rm yr}$ which is very close to the bound based on pulsars.
-There are many other testable predictions in cosmology in addition to the Hubble tension and those mentioned above.
-A varying $G_{N}$ may shift CMB acoustic peaks due to modified expansion at $z\sim 1100$, which is testable with Planck or the Simons Observatory data.
-The minimal scale factor $a_{\rm min}\sim 10^{-22}$ in eq.
+$\bullet$   The minimal scale factor $a_{\rm min}\sim 10^{-22}$ in eq.
 (58) could produce a primordial gravitational wave spectrum detectable by LIGO or eventually the space based interferometer LISA.
@@
 In fact, based on (30), for a weaker $G_{N}$, black holes are smaller for a given mass.
-Very recently, supermassive black holes, with mass on the order of 40 million solar masses have been discovered using the James Webb Space Telescope JWST which were formed at an earlier time than expected, and this may be an interesting topic for further study in the context of a varying $G_{N}$.
-$\bullet$   As stated in the Introduction, although the formula (3) is well-motivated by the work ALrhovac1 ; ALrhovac2 , the QFT underlying this formula for $\rho_{\rm vac}$} has not been precisely identified, in particular the zeron with mass $m_{\rm z}$ has not yet been identified as a known particle, though Majorana fermions are viable candidates.
-This underlying QFT is incomplete in the UV since it relies on the RG beta function (35) for a marginally irrelevant coupling, where the incompleteness is manifested as a Landau pole at very high energies.
-However we showed that the time evolution of the scale factor $a(t)$ smoothly traverses this pole through a swing rather than a bang.
-The UV completeness issue here is certainly more tractable than the UV incompleteness of quantum gravity, since it is a UV issue of QFT in flat Minkowski space where such problems are better understood.
-Landau poles are commonplace in elementary particle physics and condensed matter physics, since they are generic to marginally irrelevant perturbations.
-See for instance LandauPole1 ; LandauPole2 ; LandauPole3 and references therein.
-Although QED itself is not relevant to this article, let us mention that it also is not UV complete due to a Landau pole, however in this context it occurs at much higher energies than our proposed $z_{\rm max}$, and is thus commonly viewed as a purely academic issue, and is furthermore complicated by chiral symmetry breaking.
-(See footnote 12).
+Very recently, supermassive black holes, with mass on the order of 40 million solar masses have been discovered using the James Webb Space Telescope JWST , which were formed at an earlier time and smaller than expected, and this may be an interesting topic for further study in the context of the varying $G_{N}$ that we specifically proposed.
+$\bullet$   As stated in the Introduction, although the formula (2) is well-motivated by the works ALrhovac1 ; ALrhovac2 , we left aside the issue of identifying the particle with mass $m_{\rm z}$ underlying this formula for $\rho_{\rm vac}$} since we were able to make progress without doing so, and the particle physics involved is beyond the original scope of this article.
+Nevertheless, we can constrain some of its properties.
+We were able to constrain its mass from the observed value of $\rho_{\rm vac}$} in (5), and pointed out that this is consistent with Majorana neutrinos as viable candidates, but were not able to make a strong case for this without a more complete theory of the origin of neutrino masses.
+Based on our comparison of coupling constants with QED, it could be that the particle with mass $m_{\rm z}$ is coupled to a hidden U(1) gauge theory where $\mathfrak{g}$ is a marginally irrelevant fine-structure constant.
+We assumed that our formula for $\rho_{\rm vac}$} in (2) is valid up to $z_{\rm max}$, which would seem to imply that such a particle does not obtain its mass $m_{\rm z}$ from the Higgs mechanism since electro-weak symmetry breaking occurs at a significantly lower scale of about $160$ GeV, which corresponds to $z\approx 6\times 10^{14}$.
+In our current understanding of the Standard Model of particle physics, neutrino masses do not arise from the Higgs mechanism since the Standard Model does not have right-handed neutrinos to pair up with the known left-handed neutrinos to provide a mass.
+Based on this, a massive Majorana neutrino is the most promising candidate for the $m_{\rm z}$-particle, however other possibilities where this is an entirely new particle are certainly not ruled out.
+This underlying QFT involving $m_{\rm z}$ is incomplete in the UV since it relies on the RG beta function (47) for a marginally irrelevant coupling, where the incompleteness is manifested as a Landau pole at very high energies.
+However we showed that the time evolution of the scale factor $a(t)$ smoothly traverses this Landau pole through a swing rather than a bang.
+The UV completeness issue here is certainly more tractable than the UV incompleteness of quantum gravity, since it is a UV issue of QFT in flat Minkowski space where such problems are much better understood.
+$\bullet$   In Section VIID we described some not so well-known small scale bench-top experiments on the temperature variation of Newton’s constant which provided some positive support for our model.
+It is potentially exciting that our new parameter $\widehat{b}$ for cosmology can in principle be measured by these completely different kinds of experiments, and this justifies reproducing these results with more modern experimental techniques.
+IX Acknowledgements
+We wish to thank Maria Dainotti, Giovanni Montani, and Biagio de Simone of the NAOJ collaboration for pointing out their recent articles and discussions of their results shortly after the first version of this article appeared.
+We also benefitted from subsequent discussions with Seamus Davis and Eanna Flanagan.